### v1

# RWKV-UNet: Improving UNet with Long-Range Cooperation for Effective Medical Image Segmentation

Abstract
In recent years, there have been significant advancements in deep learning for medical image analysis, especially with convolutional neural networks (CNNs) and transformer models. However, CNNs face limitations in capturing long-range dependencies while transformers suffer high computational complexities. To address this, we propose RWKV-UNet, a novel model that integrates the RWKV (Receptance Weighted Key Value) structure into the U-Net architecture. This integration enhances the model’s ability to capture long-range dependencies and improve contextual understanding, which is crucial for accurate medical image segmentation. We build a strong encoder with developed inverted residual RWKV (IR-RWKV) blocks combining CNNs and RWKVs. We also propose a Cross-Channel Mix (CCM) module to improve skip connections with multi-scale feature fusion, achieving global channel information integration. Experiments on benchmark datasets, including Synapse, ACDC, BUSI, CVC-ClinicDB, CVC-ColonDB, Kvasir-SEG, ISIC 2017 and GLAS show that the RWKV-UNet achieves state-of-the-art performance on various types of medical image segmentation. Additionally, smaller variants, RWKV-UNet-S and RWKV-UNet-T, balance accuracy and computational efficiency, making them suitable for broader clinical applications.
1 Introduction
Computer-aided medical image analysis plays a crucial role in modern healthcare. Automated image segmentation is a critical component of this process, as it divides complex images into distinct regions, allowing specific tissues, organs, or lesions to be more clearly visible, improving diagnostic accuracy, and facilitating the development of personalized treatment plans. The rise of deep learning [28] has greatly advanced image segmentation by automatically learning intricate patterns within images with deep neural networks, improving the precision and efficiency of the analysis.
Current medical image segmentation approaches mainly rely on fully convolutional neural networks (FCNs) characterized by a U-shaped architecture. U-Net [46] is specifically designed for image segmentation tasks, particularly in the medical imaging domain. Its distinctive features include an encoder-decoder structure that enables effective downsampling and upsampling of images, along with skip connections. This design allows U-Net to capture both high-level semantic features and fine-grained details, resulting in improved segmentation accuracy. The architecture’s success has led to the development of numerous variants, such as Attention U-Net [40], U-Net++[61], U-Net3+ [22], which have further enhanced its performance and expanded its applicability.
While CNNs excel at capturing local features through fixed-size kernels, they often struggle with long-range dependencies needed for understanding complex structures. Transformers, with their self-attention mechanism, capture global context and intricate patterns across entire images. Combining these architectures allows U-Net to retain fine detail, as seen in models like TransUNet [9], MT-UNet [53], and HiFormer[21]. However, transformers have $O(N^{2})$ complexity in self-attention, which makes them computationally intensive for high-resolution scans. The trade-off between accuracy and efficiency is important in clinical scenarios, thus reducing computational complexity without significantly compromising accuracy is highly desirable.
Recently, sequence modeling methods based on linear attention mechanisms, such as Mamba  [17] and RWKV  [41], have attracted significant attention and been applied to computer vision  [62, 16] and medical image segmentation tasks  [47, 38, 33]. These models excel at capturing long-range dependencies with high accuracy, effectively preserving important contextual information across extensive image regions. In addition, their linear computational complexity greatly improves efficiency, making them suitable for large-scale or high-resolution images while reducing processing demands.
Inspired by Vision-RWKV [47], we propose a novel RWKV-UNet, integrating convolutional layers with the spatial mix for medical image segmentation. The motivation behind this design is to take advantage of RWKV’s ability to capture long-range dependencies and combine it with the strengths of convolution localization, in order to achieve more precise and context-sensitive segmentation results. To evaluate our model’s effectiveness, we conducted experiments on benchmark datasets including Synapse  [27], ACDC  [5], BUSI  [1], CVC-ClinicDB  [4], CVC-ColonDB  [3], Kvasir-SEG  [25], ISIC 2017  [13] and GLAS  [49], where it achieves state-of-the-art (SOTA) performance, highlighting its potential as a robust tool for medical image analysis.
In summary, our contributions are threefold.
  • We propose a novel U-Net architecture based on a combination of RWKV and convolutional layers. An encoder based on the inverted residual RWKV (IR-RWKV) module is designed, effectively combining global and local information, with enhanced feature extraction capabilities through pre-training.
  • A Cross-Channel Mix module is designed to improve the skip connections in U-Net, facilitating multi-scale information fusion and enhancing the model’s feature representation across different scales.
  • Experiments conducted on datasets of different modalities show that the model achieves SOTA performance, demonstrating strong segmentation capabilities. We also propose relatively smaller models, RWKV-UNet-S and RWKV-UNet-T, which strike a balance between performance and computational efficiency, making them suitable for a wider range of scenarios.
2 Related Work
2.1 UNet Variants
U-Net is a deep learning architecture designed for image segmentation, originally developed for biomedical images. Its symmetrical U-shaped structure consists of an encoder and a decoder, where the encoder progressively extracts features and reduces spatial dimensions, while the decoder gradually restores spatial information and generates high-resolution outputs. Skip connections are used to combine feature maps from the encoder layers with those from the decoder layers, ensuring the preservation of details. Variations in U-Net improve its performance through various means, including the use of deeper encoders such as ResNet  [59], the introduction of gating mechanisms to focus on important features  [40], improvements in skip connection formats  [61, 22, 64, 42] to recover more spatial details, as well as adaptations for 3D images [12, 39]. Some studies have explored combining MLP/KAN with UNet to achieve long-range dependencies  [50, 35, 30]. These improvements allow U-Net to demonstrate higher accuracy and adaptability in various image segmentation tasks.
Attention-based Improvements.
Attention-based U-Net has several advantages, primarily its ability to capture global context information, which enhances the model’s understanding of long-range dependencies. Furthermore, the self-attention mechanism of Transformers allows the model to focus more effectively on key information. Variants of attention-based U-Net include pure Transformer models  [19, 23, 2, 8], which rely entirely on self-attention mechanisms for feature extraction and image segmentation. Another category consists of hybrid architectures that combine Transformers with CNNs  [9, 7, 53, 57, 52, 55, 21, 26], which leverage the local feature extraction capabilities of CNNs along with the global information processing strengths of Transformers to better capture image details and structural information. Recently, Mamba  [17]’s success in NLP and vision tasks  [62] has drawn significant attention to SSM, which can be considered a form of linear attention. SSM-based UNet variants  [47, 38, 54, 34, 33] capture long-range dependencies effectively and offer linear computational complexity, enhancing accuracy and efficiency in image segmentation tasks.
2.2 Receptance Weighted Key Value (RWKV)
Receptance Weighted Key Value (RWKV)  [41] was initially developed for natural language processing tasks. It combines the linear complexity of RNNs with the parallel processing advantages of Transformers, resulting in excellent performance, constant memory usage, and consistent inference generation speed. Building on this, Vision-RWKV (VRWKV)  [16] introduces RWKV to computer vision tasks, achieving success in image classification, object detection, and semantic segmentation. It efficiently handles sparse inputs and robust global processing and allows for high-resolution image processing without the need for window operations. Related efforts have also emerged in medical image analysis tasks. Restore-RWKV  [56] is an RWKV-based model for medical image restoration, enhancing the efficiency and effectiveness of medical image restoration through the introduction of the linear complexity Re-WKV attention mechanism and the omnidirectional token shift (Omni-Shift) layer. BSBP-RWKV [60, 60] achieves superior accuracy in medical image segmentation by combining the noise reduction advantages of Perona-Malik diffusion with the efficient RWKV structure.
3 Methodology: RWKV-UNet
3.1 Overall Architecture
Figure 1: Overall architecture of the proposed RWKV-UNet. (a) the encoder constructed by stacked IR blocks and stacked IR-RWKV blocks, (b) Cross-Channel Mix (CCM) Module, (c) the decoder constructed by point convolution layers and depthwise convolution layers,(d) the Inverted Residual (IR) block, (e) the Inverted Residual RWKV (IR-RWKV) block, (f) the architecture of the decoder block, containing a point-convolution layer and a $\times 9$ 9 DW-Conv layer, following a point operation layer and an upsampling operation.
The overall architecture of the proposed RWKV-UNet is presented in Figure 1. RWKV-UNet consists of an encoder, a decoder and skip connections with a Cross-Channel Mix (CCM) Module.
3.2 Effective Encoder in RWKV-UNet
Inverted Residual RWKV (IR-RWKV) Block.
We employ the spatial-mix module in Vision RWKV and Depth-Wise Convolution (DW-Conv) with a skip connection to combine global and local dependency, and trade-off model cost and accuracy. The process of expanding and reducing dimensions enhances feature representation by capturing richer details while preventing information bottlenecks. Given a feature map $X\in\mathbb{R}^{C_{\mathrm{in}}\times H\times W}$, the processing flow through the IR-RWKV module is as follows:
Normalize and project the input $X$ to an intermediate dimension $C_{mid}$ using a $\times 1$ 1 Convolution layer:
$I_{1}=\text{LayerNorm}(\text{Conv}_{1\times 1}(X)),$ (1)
where $I_{1}\in\mathbb{R}^{C_{\mathrm{mid}}\times H\times W}$. $C_{\mathrm{mid}}$ should be greater than $C_{\mathrm{in}}$ to achieve dimension expansion.
Divide the feature map into patches of size $\times 1$ 1. Then the Spatial Mix in VRWKV is applied:
$I_{2}=\text{Unfolding}(I_{1}),$ (2)
$I_{3}=\text{SpatialMix}(\text{LayerNorm}(I_{2}))+I_{2},$ (3)
where $I_{2}\in\mathbb{R}^{C_{\mathrm{mid}}\times N}$, $I_{3}\in\mathbb{R}^{C_{\mathrm{mid}}\times N}$.
Convert the feature sequence back into a 2D feature map $I_{4}=\text{Folding}(I_{3})\in\mathbb{R}^{C_{\mathrm{mid}}\times H\times W}$, and use a $\times 5$ 5 DW-Conv layer for local feature aggregation:
$I_{5}=\text{DW-Conv}(I_{4})+I_{4},$ (4)
where $I_{5}\in\mathbb{R}^{C_{\mathrm{mid}}\times H^{\prime}\times W^{\prime}}$. $H^{\prime}\times W^{\prime}$ is determined by the stride of DW-Conv.
Finally, project $I_{5}$ to the output dimension and add a global skip connection:
$F=\text{Conv}_{1\times 1}(I_{5})+X,$ (5)
where $F\in\mathbb{R}^{C_{\mathrm{out}}\times H^{\prime}\times W^{\prime}}$.
Inverted Residual (IR) Block.
An IR block contains a point convolution layer, a DW-Conv layer and a point convolution layer with local and global residual skips, removing the spatial mix as well as the unfolding and folding processes from the IR-RWKV.
Effective Encoder with Stacked IRs and IR-RWKVs.
We construct the RWKV-UNet encoder by stacking the IR and IR-RWKV blocks. The encoder comprises a stem stage followed by four stages. The first blocks in stages \@slowromancapi@, \@slowromancapii@, \@slowromancapiii@ and \@slowromancapiv@ do not have residual connections and use a DW-Conv with a stride of 2 to achieve downsampling. After the first block, Stages \@slowromancapi@ and \@slowromancapii@ are composed of stacked IR blocks without Spatial Mix. In Stages \@slowromancapiii@ and \@slowromancapiv@, a series of IR-RWKV blocks follow the first block.
Scale Up.
As shown in Table 1, we implement three different sizes of encoders with different numbers of IR and IR-RWKV blocks in each stage (depths), different embedding dimensions and expansion ratios for $C_{\mathrm{mid}}$ : Encoder-Tiny (Enc-T), Encoder-Small (Enc-S), and Encoder-Base (Enc-B), allowing flexibility to scale up the model according to different resource constraints and application needs.
Pre-training Manner.
We pre-train the encoders of different sizes on ImageNet-1K  [14] with a batch size of 1024 for 300 epochs using the AdamW  [36] optimizer. These pre-trained weights of different sizes will be used in medical segmentation tasks, enabling improved feature extraction and faster convergence.
Table 1: RWKV-UNet Encoder Configurations Across Stages
Stage Parameter Enc-T Enc-S Enc-B
Stem Dimension 24 24 24
Stage \@slowromancapi@ Depth 2 3 3
Dimension 32 32 48
Expansion 2.0 2.0 2.0
Spatial Mix
Stage \@slowromancapii@ Depth 2 3 3
Dimension 48 64 72
Expansion 2.5 2.5 2.5
Spatial Mix
Stage \@slowromancapiii@ Depth 4 6 6
Dimension 96 128 144
Expansion 3.0 3.0 4.0
Spatial Mix
Stage \@slowromancapiv@ Depth 2 3 3
Dimension 160 192 240
Expansion 3.5 4.0 4.0
Spatial Mix
Parameters 2.6M 9.4M 16.7M
FLOPs 1.9G 4.8G 9.0G
Table 2: Comparison results of models with different attention mechanisms on Synapse dataset.
Attention Type Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
No Attention 18.06 77.05 04.55G
Multi-Head Self-attention [51] 16.95 77.60 14.89G
Focused Linear Attention [18] 18.45 75.86 11.69G
Bi-Mamba [62] 13.13 77.86 14.53G
Spatial-Mix [47] 14.03 78.68 11.22G
Comparison with other attention types
We conduct experiments by replacing the spatial mix in the encoder of RWKV-UNet with other attention mechanisms, such as self-attention  [51], focused linear attention  [18] and Bi-Mamba used in Vision Mamba  [62], as well as by removing this module entirely. The total training epochs are 60. The results shown in Table 2 indicate that the model with Spatial-Mix achieves the best DSC with the lowest computational cost compared to other attention mechanisms, showcasing its efficiency and effectiveness. And all kinds of attention mechanisms can lead to improvement compared to pure CNN structure except focused linear attention.
3.3 Decoder Design
We design our CNN-based decoder block with a point-convolution layer and a $\times 9$ 9 DW-Conv layer, following a point operation layer and an upsampling operation. Consider the input feature map $X\in\mathbb{R}^{C_{\text{in }}\times H\times W}$; in depthwise convolution, the number of input channels is $C_{\mathrm{in}}$, and the second point convolution maps the $C_{\text{in }}$ channels to $C_{\text{out }}$. The computational complexity of a standard convolution is:
$H\times W\times C_{\text{in }}\times\left(C_{\text{out }}\times K^{2}\right).$ (6)
The computational complexity of pointwise convolution + depthwise convolution + pointwise convolution is:
$H\times W\times C_{\text{in }}\times\left(C_{\text{in }}+K^{2}+C_{\text{out }}\right),$ (7)
where $K$ is the kernel size. The latter method is more computationally efficient for the use of larger kernels, as the complexity is proportional to $C_{\text{in }}\times K^{2}$ rather than $C_{\text{in }}\times C_{\text{out }}\times K^{2}$.
3.4 Cross-Channel Mix for Multi-scale Fusion
Inspired by the Channel Mix in VRWKV, we propose Cross-Channel Mix (CCM), a module that can effectively extract channel information across multi-scale encoder features. By capturing the rich global context along channels, CCM further enhances the extraction of long-range contextual information.
Define output feature map of Stage \@slowromancapi@, Stage \@slowromancapii@ and Stage \@slowromancapiii@ of the designed encoder are $F_{1},F_{2}$, and $F_{3}$, which has different Spatial dimensions and channel counts: $F_{1}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}},\quad F_{2}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}},\quad F_{3}\in\mathbb{R}^{C_{3}\times H_{3}\times W_{3}}$. ($H_{1}$ >$H_{2}$>$H_{3}$, $W_{1}$ >$W_{2}$ >$W_{3}$ and $C_{3}$ >$C_{2}$ >$C_{1}$).
We reshape and map smaller features to match the largest feature’s size and dimension:
$\begin{gathered}\tilde{F}_{i}=\operatorname{Conv}_{C_{i}}\left(\operatorname{Upsample}\left(F_{i}\right)\right),\quad i=2,3\\ \tilde{F}_{1}=\operatorname{Conv}_{C_{1}}\left(F_{1}\right),\end{gathered}$ (8)
where $\tilde{F}_{1}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$, $\tilde{F}_{2}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$, and $F_{3}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$.
We concatenate the adjusted feature maps along the channel dimension to produce a combined feature map:
$F_{\text{cat}}=\text{Concat}(\tilde{F}_{1},\tilde{F}_{2},\tilde{F}_{3}),$ (9)
where $F_{\text{cat}}\in\mathbb{R}^{3C_{1}\times H_{1}\times W_{1}}$.
Divide the feature map into patches: $F_{\text{unf}}=\text{Unfolding}(F_{\text{cat}})\in\mathbb{R}^{3C_{1}\times N}$. Then apply the Channel Mix in VRWKV, performing multi-scale global feature fusion in the channel dimension:
$F_{\text{mix}}=\text{ChannelMix}(\text{LayerNorm}(F_{\text{unf}}))+F_{\text{unf}},$ (10)
where $F_{\text{mix}}\in\mathbb{R}^{3C_{1}\times N_{1}}$. Project it back to 2D feature, $F_{\text{fold}}=\text{Folding}(F_{\text{mix}})\in\mathbb{R}^{3C_{1}\times H_{1}\times W_{1}}$, and $N_{1}=H_{1}\times W_{1}$.
[…]
$F_{i}^{\prime}=\operatorname{Conv}_{C_{i}}\left(\operatorname{Reshape}\left(F_{\text{fold }}^{(i)}\right)\right),\quad i=1,2,3$ (11)
where $F_{1}^{\prime}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$ , $F_{2}^{\prime}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$ and $F_{3}^{\prime}\in\mathbb{R}^{C_{3}\times H_{3}\times W_{3}}$. $F_{1}^{\prime}$,$F_{2}^{\prime}$ and $F_{3}^{\prime}$ will be concatenated with the output of the decoder Stage \@slowromancapiii@, Stage \@slowromancapii@ and Stage \@slowromancapi@.
4 Experiments
Figure 2: Average DSC vs. #Params for different methods on Synapse multi-organ segmentation dataset.
Figure 3: Average DSC vs. #FLOPs for different methods on Synapse multi-organ segmentation dataset.
[…]
Methods Average Aotra Gallbladder Kidney (Left) Kidney (Right) Liver Pancreas Spleen Stomach
HD95 $\downarrow$ DSC $\uparrow$
U-Net  [46] 39.70 76.85 89.07 69.72 77.77 68.60 93.43 53.98 86.67 75.58
R50 U-Net  [20, 46] 36.87 74.68 87.74 63.66 80.60 78.19 93.74 56.90 85.87 74.16
R50 Att-Unet  [20, 40] 36.97 75.57 55.92 63.91 79.20 72.71 93.56 49.37 87.19 74.95
ViT  [15] + CUP [9] 36.11 67.86 70.19 45.10 74.70 67.40 91.32 42.00 81.75 70.44
R50-ViT  [20, 15] + CUP [9] 32.87 71.29 73.73 55.13 75.80 72.20 91.51 45.99 81.99 73.95
TransUNet  [9] 31.69 77.48 87.23 63.16 81.87 77.02 94.08 55.86 85.08 75.62
SwinUNet  [6] 21.55 79.13 85.47 66.53 83.28 79.61 94.29 56.58 90.66 76.60
TransClaw U-Net  [7] 26.38 78.09 85.87 61.38 84.83 79.36 94.28 57.65 87.74 73.55
MT-UNet  [53] 26.59 78.59 87.92 64.99 81.47 77.29 93.06 59.46 87.75 76.81
UCTransNet  [52] 26.75 78.23 - - - - - - - -
MISSFormer  [23] 18.20 81.96 86.99 68.65 85.21 82.00 94.41 65.67 91.92 80.81
CASTformer [57] 22.73 82.55 89.05 67.48 86.05 82.17 95.61 67.49 91.00 81.55
TransDeepLab [2] 21.25 80.16 86.04 69.16 84.08 79.88 93.53 61.19 89.00 78.40
LeVit-Unet-384  [55] 16.84 78.53 87.33 62.23 84.61 80.25 93.11 59.07 88.86 72.76
CTC-Net  [58] - 78.41 86.46 63.53 83.71 80.79 93.78 59.73 86.87 72.39
MS-UNet  [8] 18.97 80.44 85.80 69.40 85.86 81.66 94.24 57.66 90.53 78.33
HiFormer-L  [21] 19.14 80.69 87.03 68.61 84.23 78.37 94.07 60.77 90.44 82.03
PVT-CASCADE  [44] 20.23 81.06 83.01 70.59 82.23 80.37 94.08 64.43 90.10 83.69
TransCASCADE  [44] 17.34 82.68 86.63 68.48 87.66 84.56 94.43 65.33 90.79 83.52
AiA-UNet  [43] - 82.88 87.77 74.33 84.16 82.39 94.98 65.65 92.77 80.99
MetaSeg-B  [26] - 82.78 - - - - - - - -
VM-UNet  [47] 19.21 81.08 86.40 69.41 86.16 82.76 94.17 58.80 89.51 81.40
PVT-EMCAD-B2  [45] 15.68 83.63 88.14 68.87 88.08 84.10 95.26 68.51 92.17 83.92
SelfReg-UNet  [63, 46] - 80.34 88.74 71.78 85.32 80.71 93.80 62.22 84.78 75.39
SelfReg-SwinUnet  [63, 6] - 80.54 86.07 69.65 85.12 82.58 94.18 61.08 87.42 78.22
RWKV-UNet-T (ours) 17.83 81.58 89.28 64.18 86.03 83.76 95.09 63.50 90.36 80.44
RWKV-UNet-S (ours) 21.16 82.20 89.36 67.60 84.11 79.26 95.60 68.86 91.58 81.21
RWKV-UNet (ours) 15.70 84.02 89.53 68.94 87.63 84.07 95.57 69.38 90.95 86.09
Table 4: Comparison results on ACDC dataset. The evaluation metric is DSC in (%).
Methods Average RV Myo LV
R50 U-Net  [20, 46] 87.55 87.10 80.63 94.92
ViT  [15] + CUP [9] 81.45 81.46 70.71 92.18
R50-ViT  [20, 15] + CUP [9] 87.57 86.07 81.88 94.75
R50 Att-UNet  [20, 40] 86.75 87.58 79.20 93.47
TransUNet  [9] 89.71 88.86 84.54 95.73
SwinUNet  [6] 90.00 88.55 85.62 95.83
MT-UNet  [53] 90.43 86.64 89.04 95.62
MS-UNet  [8] 87.74 85.31 84.09 93.82
MISSFormer  [23] 90.86 89.55 88.04 94.99
CTC-Net  [58] 90.77 90.09 85.52 96.72
LeViT-UNet-384s  [55] 90.32 89.55 87.64 93.76
PVT-CASCADE  [44] 90.45 87.20 88.96 95.19
TransCASCADE  [44] 91.63 89.14 90.25 95.50
PVT-EMCAD-B2  [45] 92.12 90.65 89.68 96.02
SelfReg-UNet  [63, 46] 91.43 88.92 89.49 95.88
SelfReg-SwinUnet [63, 6] 91.49 89.49 89.27 95.70
RWKV-UNet-T (ours) 90.92 89.73 87.18 95.85
RWKV-UNet-S (ours) 91.64 89.77 88.48 96.65
RWKV-UNet (ours) 92.17 90.86 88.72 96.92
Table 5: Comparison results on BUSI, CVC-ClinicDB, CVC-ColonDB, Kvasir-SEG, ISIC 2017 and GLAS datasets. The evaluation metric are DSC in (%). Results with error bars are averaged over three runs with three different splits.
Methods BUSI CVC-ClinicDB CVC-ColonDB Kvasir-SEG ISIC 2017 GLAS Params. FLOPs
U-Net [46] 74.90$\pm$1.15 89.22$\pm$4.07 83.43$\pm$5.24 84.94$\pm$2.60 82.94 88.56$\pm$0.63 07.8M 12.15G
Att-UNet [40] 78.14$\pm$1.25 90.86$\pm$4.09 84.75$\pm$6.55 86.21$\pm$2.15 83.96 89.14$\pm$0.54 08.7M 16.79G
UNet++ [61] 75.74$\pm$1.70 90.07$\pm$5.09 84.75$\pm$3.99 85.86$\pm$1.52 82.27 88.93$\pm$0.67 09.2M 34.77G
TransUNet [9] 79.35$\pm$2.45 93.91$\pm$2.28 90.72$\pm$3.92 91.34$\pm$0.01 86.09 91.58$\pm$0.15 00.1G 38.54G
FCBFormer [48] 80.66$\pm$0.28 91.15$\pm$3.50 85.23$\pm$5.21 91.09$\pm$0.96 84.56 91.13$\pm$0.25 53.0M 41.60G
PVT-CASCADE  [44] 78.30$\pm$3.17 93.45$\pm$1.67 90.97$\pm$2.11 91.40$\pm$1.29 84.77 91.85$\pm$0.39 35.3M 08.40G
Rolling-UNet-M  [31] 77.49$\pm$2.42 91.91$\pm$2.56 85.20$\pm$5.85 87.41$\pm$1.78 84.42 88.12$\pm$0.66 07.1M 08.31G
PVT-EMCAD-B2  [45] 79.22$\pm$1.07 93.39$\pm$1.64 91.27$\pm$1.84 90.37$\pm$1.10 84.05 91.77$\pm$0.09 26.8M 05.85G
U-KAN  [30] 76.95$\pm$1.67 89.88$\pm$3.21 84.66$\pm$5.33 85.78$\pm$1.68 83.46 87.42$\pm$0.28 25.4M 10.59G
RWKV-UNet-T (ours) 79.47$\pm$1.35 95.05$\pm$0.67 89.56$\pm$5.26 91.26$\pm$2.13 85.59 91.96$\pm$0.31 03.3M 03.78G
RWKV-UNet-S (ours) 80.92$\pm$1.74 95.58$\pm$0.69 91.30$\pm$3.21 91.78$\pm$1.24 86.38 92.11$\pm$0.16 09.9M 07.75G
RWKV-UNet (ours) 81.92$\pm$0.82 95.26$\pm$0.70 92.27$\pm$2.48 91.26$\pm$1.27 85.32 92.35$\pm$0.07 17.4M 14.64G
4.1 Experimental Setup
Datasets and Metrics.
Experiments on conducted on Synapse [27] for multi-organ Segmentation in CT images, ACDC [5] for cardiac segmentation in MRI images, BUSI [1] for breast tumor segmentation in ultrasound images, CVC-ClinicDB [4], CVC-ColonDB [3], Kvasir-SEG [25] for poly segmentation, ISIC 2017 [13] for skin lesion segmentation and GLAS  [49] for gland segmentation. The average Dice Similarity Coefficient (DSC) and average 95% Hausdorff Distance (HD95) are used as evaluation metrics. Specific details for datasets can be seen in Appendix.
Comparison Methods.
We use U-Net  [46], R50 U-Net  [20, 46], R50 Att-Unet  [20, 40], ViT  [15] + CUP  [9], R50-ViT  [20, 15] + CUP  [9], TransUNet  [9], SwinUNet  [6], TransClaw U-Net  [7], MT-UNet  [53], UCTransNet  [52], MISSFormer  [23], CASTformer  [57], TransDeepLab  [2], CTC–Net  [58], MS-UNet  [8], LeVit-Unet-384  [55], HiFormer-L  [21], PVT-CASCADE  [44], TransCASCADE  [44], AiA-UNet  [43], MetaSeg-B  [26], VM-UNet  [47],PVT-EMCAD-B2  [45], SelfReg-UNet  [63, 46], SelfReg-SwinUnet  [63, 6], FCBFormer [48], Rolling-UNet-M  [31], U-KAN  [30]as baselines.
Implementation Details.
All experiments are all performed on PG500-216(V-100) with 32 GB memory. The resolution of the input images is resized to 224$\times$224 for Synapse and ACDC, and 256$\times$256 for binary segmentation tasks. The total training epochs for Synapse are 30, for ACDC are 150, and for binary segmentation are 300. The batch size is 24 for Synapse and ACDC, and 8 for binary segmentation tasks. The initial learning rate is 1e-3 for Synapse, 5e-4 for ACDC, and 1e-4 for binary segmentation experiments. The minimum learning rates for Synapse and ACDC are 0 and for binary segmentation are 1e-5. The AdamW  [36] optimizer and CosineAnnealingLR  [37] scheduler are used. Test-time augmentation is used for Synapse and ACDC. The baseline results on binary segmentation tasks are run by ourselves. The loss function used is a mixed loss combining cross entropy (CE) loss and dice loss  [39]:
$\mathcal{L}=\alpha CE(\hat{y},y)+\beta Dice(\hat{y},y),$ (12)
where $\alpha$ and $\beta$ are 0.5 and 0.5 for Synapse, 0.3 and 0.7 for ACDC, 0.5 and 1 for binary segmentation. Experiments for PVT-CASCADE  [44] and PVT-EMCAD-B2  [45] use deep supervision strategy  [29].
Figure 4: A qualitative comparison with previous SOTA methods on the Synapse dataset. The visual results demonstrate that our method achieves more accurate segmentation, especially in difficult tasks like pancreas segmentation.
4.2 Comparison with State-of-the-arts
Multi-organ Segmentation.
Table 3 shows that RWKV-UNet excels in abdomen organ segmentation on the Synapse multi-organ dataset, achieving the highest average DICE score of 84.02% and surpassing all SOTA CNN- and transformer-based methods. It also achieves 15.70 in HD95, only slightly lower than PVT-EMCAD-B2, demonstrating a strong ability to localize the organ boundary. Its performance on the most challenging task, pancreas segmentation, is the best, indicating its effectiveness in handling sample imbalance issues. Smaller models, RWKV-UNet-T and RWKV-UNet-S, also achieve remarkable results.
Cardiac Segmentation.
Table 4 shows that our model outperforms all SOTA models in the ACDC MRI cardiac segmentation task, achieving the best results in both RV and LV categories. The smaller models, RWKV-UNet-T and RWKV-UNet-S, also perform well.
Binary Segmentation.
Table 5 shows that RWKV-UNet series show remarkable performance in the tasks of breast lesion, poly, skin lesion and grand segmentation, with much smaller parameters and computation load than TransUNet and FCBFormer. Meanwhile, it is observed that in some datasets the smaller model RWKV-UNet-S perform better.
4.3 Ablation and Analysis
Effect of Pre-training for Encoder.
Whether to use pre-trained weights of RWKV-UNet encoder results in Table 6. Experimental results indicate that pre-training is essential because it significantly improves feature extraction by enabling the encoder to capture intricate patterns and hierarchical information. Notably, without pre-training, HD95 is higher. But DSC is a more important metric, providing a reliable assessment of segmentation quality directly, while HD95 can be overly sensitive to outliers.
Table 6: Comparison results on Synapse dataset of whether to use pre-trained weights on ImageNet-1k or not.
Pre-training Training Epochs Average
HD95 $\downarrow$ DSC $\uparrow$
w/o 30 13.94 76.49
w/o 60 14.03 78.68
w/ 30 15.70 84.02
Table 7: Comparison results of different decoder designs for RWKV-UNet on Synapse dataset. $c\rightarrow c\rightarrow d$ denotes $\operatorname{Conv}_{1\times 1}\rightarrow\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv and $c\rightarrow d\rightarrow c$ denotes $\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv $\rightarrow\operatorname{Conv}_{1\times 1}.$
Decoder Kernel Size Average Flops
HD95 $\downarrow$ DSC $\uparrow$
Conv + upsample 3 24.95 81.77 11.38G
$c\rightarrow c\rightarrow d$
+ upsample
3 23.44 82.10 11.06G
5 25.53 81.84 11.07G
7 20.78 81.57 11.08G
9 18.07 83.76 11.10G
$c\rightarrow d\rightarrow c$
+ upsample
3 21.68 83.33 11.08G
5 22.26 83.74 11.11G
7 22.90 82.54 11.16G
9 15.70 84.02 11.22G
Decoder Design.
In Tabel 7, we examine various sizes of convolution kernels, finding that the use of large kernels, such as $\times 9$ 9, in the decoder significantly enhances the segmentation performance. This improvement is likely due to larger kernels capturing broader contextual information and preserving more spatial details, which enhances segmentation accuracy in medical images. We also compare different ways to construct the decoder block and the results show that $\operatorname{Conv}_{1\times 1}\rightarrow\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv performs better than $\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv $\rightarrow\operatorname{Conv}_{1\times 1}$ and the common convolution layer, showing that DW-Conv operating on larger dimensions will help the decoding process.
Channel Mix.
In VRWKV [16], a Channel Mix module is used after the Spatial Mix. In RWKV-UNet’s encoder, IR-RWKV blocks are designed without using Channel Mix as the pointwise convolution in the output layer performs both channel mix and feed-forward operations. We compare the results on the Synapse dataset with and without applying channel mix after spatial mix. All experiments are performed without pre-training, with a total of 60 epochs. The results are shown in Table 8, demonstrating that the channel mix is not essential for maintaining performance.
Table 8: Comparison of results on the Synapse dataset: effect of applying Channel Mix after Spatial Mix in the IR-RWKV blocks
Channel Mix Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
w/ 15.29 78.76 19.49G
w/o (ours) 14.03 78.68 11.22G
Skip Connections.
Comparison results of different skip connection designs for RWKV-UNet on the Synapse dataset are shown in Table 9. Experiments analyze variations in the number of skip connections and the inclusion of the CCM block. We prioritize retaining shallow skip connections because shallow layers tend to experience greater information loss. The results show that, like other UNet’s variants, adequate skip connections are essential for RWKV-UNet’s performance. Furthermore, the results show that the CCM module can significantly improve the model segmentation results. Even if only one or two inputs from the encoder(when only one input, the CCM block degrades to a ”channel mix bridge”), the CCM block can still lead to increments in DSC. Admittedly, the CCM module increases the computational load to some extent due to the aggregation of global information.
Table 9: Comparison results on different skip connections design of RWKV-UNet on Synapse dataset. Experiments are conducted on the number of skips and whether or not to use the CCM block.
Skips CCM Block Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
0 w/o 23.39 74.19 09.25 G
1 w/o 17.17 79.76 09.41 G
1 w/ 19.56 79.98 09.63 G
2 w/o 16.08 80.90 09.49 G
2 w/ 19.97 81.11 10.24 G
3 w/o 22.72 81.14 09.56 G
3 w/ 15.70 84.02 11.22 G
5 Conclusion
In this study, we introduce RWKV-UNet, a novel architecture that integrates the RWKV architecture with U-Net. By combining the strengths of convolutional networks for local feature extraction and RWKV’s ability to model global context, our model significantly improves medical image segmentation accuracy. The proposed enhancements, such as the IR-RWKV module and the CCM module, contribute to a more precise feature representation and information fusion across different scales. Experimental results on benchmark datasets demonstrate that RWKV-UNet surpasses all SOTA CNN- and transformer-based methods. Its variants (RWKV-UNet-S and RWKV-UNet-T) offer a practical balance between performance and computational efficiency. Our approach has a strong potential to advance medical image analysis, particularly in clinical settings where both accuracy and efficiency are paramount.
Limitations and Future Work. RWKV-UNet is a powerful 2D medical image segmentation model that effectively combines RWKV with convolution, but it is currently not applicable to 3D imaging. In the future, we plan to extend the model to 3D for handling volume data and to explore the potential of RWKV-based foundational models for medical imaging. We also aim to develop ultra-lightweight RWKV-based models tailored for ’bedside healthcare’ applications, preserving segmentation accuracy while enhancing adaptability and speed further.
Appendix
The supplementary material gives more specific details of our experimental settings, and presents more comprehensive analysis and results of our RWKV-UNet to facilitate the comparison of subsequent methods:
  • Sec. A provides introduction to datasets and corresponding splitting settings used in this paper.
  • Sec. B.1 provides comparison results on Synapse dataset to explore the effect of applying DW-Conv after Spatial Mix in IR-RWKV blocks.
  • Sec. B.2 provides comparison results of different methods on Synapse dataset with larger inputs.
Appendix A Introduction to datasets
1) The Synapse multi-organ segmentation dataset  [27]: 30 abdominal CT scans from the MICCAI 2015 Multi-Atlas Abdomen Labeling Challenge are used. The dataset is split into 18 training cases and 12 cases for validation following the setting in  [9].
2) The Automated Cardiac Diagnosis Challenge (ACDC)  [5]: This dataset gathezrs examinations obtained by MRI scanners. Each patient scan is manually annotated with ground truth data for the left ventricle (LV), right ventricle (RV), and myocardium (MYO). The dataset is split into 70 training cases, along with 10 cases for validation and 20 for testing following the setting in  [9].
3) Breast Ultrasound Images Dataset (BUSI) [1]: This dataset contains breast ultrasound images collected from women between 25 and 75 years old. The dataset consists of 780 pairs of images in normal, benign, and malignant, with corresponding masks. Unlike the setting in [50], we include all three classes in the dataset. The dataset is split into the training set, the validation set and the validation set with a ratio of 8:1:1 with three different random states.
4) CVC-ClinicDB  [4]:This dataset contains frames extracted from colonoscopy videos that contain several examples of polyps. The dataset consists of 612 pairs of images and corresponding masks. he dataset is split into the training set, the validation set and the testing set with a ratio of 8:1:1 with three different random states.
5) CVC-ColonDB [3]: This dataset includes 380 images, each with a resolution of 574$\times$500 pixels. These images were selected from a pool of 12,000 images obtained from 15 short colonoscopy videos, with only these 380 images being annotated. The dataset is split into the training set, the validation set and the testing set with a ratio of 8:1:1 with three different random states.
6) Kvasir-SEG  [25]: This dataset comprises 1000 polyp images and their corresponding ground truth masks with resolutions ranging from 332$\times$487 to 1920$\times$1072 pixels. he dataset is split into the training set, the validation set and the testing set with a ratio of 8:1:1 with three different random states.
7) ISIC 2017  [13]: This dataset contains skin lesions images for Skin Lesion Analysis toward Melanoma Detection Challenge at the International Symposium on Biomedical Imaging (ISBI) 2017. There are 2000 pairs of images and corresponding masks in the training set, 150 pairs in the validation set, and 600 pairs in the testing set.
8) GlaS  [49]: This dataset is designed for gland segmentation. It consists of 165 whole-slide images (WSIs), with 85 images designated for training and 80 for testing. The training dataset is split into a new training set and a validation set with a ratio of 9:1 with three different random states.
Appendix B More Ablation and Exploration Analysis
B.1 Depth-wise Convolution in IR-RWKV Blocks
The comparison results to explore the effect of DW-Conv after Spatial Mix in IR-RWKV blocks are shown in Table A1. The inclusion of DW-Conv demonstrates a notable improvement in performance metrics. The slight increase in computational cost is minimal and well-justified given the significant gains in performance. This indicates that in IR-RWKV blocks, DW-Conv plays a critical role. It provides local information that complements the long-range dependencies captured by Spatial Mix, effectively enhancing model accuracy. Without pre-training, relying solely on Spatial Mix without the inclusion of local information from DW-Conv limits the model’s ability to fully capture fine-grained spatial details. This demonstrates the necessity of integrating both components to achieve a balanced representation of local and global features.
Table A1: Comparison of results on the Synapse dataset: effect of applying DW-Conv after Spatial Mix in IR-RWKV blocks
DW-Conv Average FLOPs
HD95 $\downarrow$ DSC $\uparrow$
w/o 16.55 75.56 11.14G
w/ (ours) 14.03 78.68 11.22G
B.2 Larger Resolution Inputs
The comparison results of different methods with 512$\times$ 512 input in the Synapse dataset are shown in Table A2. The average Dice scores of other methods are from  [10]. Our results indicate that when the input resolution is increased to 512$\times$512, our model demonstrates an improvement in performance. Although the improvement is not significant, it still maintains the best performance overall. Furthermore, compared to TransUNet, the computational overhead associated with the resolution increase is less severe, showcasing the efficiency of our model at higher resolutions.
Figure A1: #FLOPs vs input resolutions. for different methods. The input channel is 1 and the output is 9 (aligned with settings on Synapse dataset).
Table A2: Comparison of different methods with 512$\times$512 resolution input on the Synapse dataset. The evaluation metric are DSC in (%). The average DSC of other methods are from  [10].
Method Average FLOPs
U-Net  [46] 81.34 -
Pyramid Attention  [32] 80.08 -
DeepLabv3+  [11] 82.50 -
UNet++  [61] 81.60 -
Attention U-Net  [40] 80.88 -
nnU-Net  [24] 82.92 -
TransUNet  [9] 84.36 168.7G
RWKV-UNet (ours)  [9] 85.52 058.6G
As shown in Figure A1, the RWKV-UNet series demonstrates advantages in handling high-resolution inputs, as shown in the FLOPs comparison. Particularly for larger resolutions, the RWKV-UNet models maintain significantly lower computational costs compared to other high-performance architectures like transUNet and MT-UNet. It demonstrates that linear attention outperforms traditional Transformers in efficiency, enabling faster and more memory-efficient processing of high-resolution inputs and long sequences.

### v2

@@
 Abstract
-In recent years, there have been significant advancements in deep learning for medical image analysis, especially with convolutional neural networks (CNNs) and transformer models.
-However, CNNs face limitations in capturing long-range dependencies while transformers suffer high computational complexities.
+In recent years, significant advancements have been made in deep learning for medical image segmentation, particularly with convolutional neural networks (CNNs) and transformer models.
+However, CNNs face limitations in capturing long-range dependencies, while transformers suffer from high computational complexity.
 To address this, we propose RWKV-UNet, a novel model that integrates the RWKV (Receptance Weighted Key Value) structure into the U-Net architecture.
-This integration enhances the model’s ability to capture long-range dependencies and improve contextual understanding, which is crucial for accurate medical image segmentation.
-We build a strong encoder with developed inverted residual RWKV (IR-RWKV) blocks combining CNNs and RWKVs.
+This integration enhances the model’s ability to capture long-range dependencies and to improve contextual understanding, which is crucial for accurate medical image segmentation.
+We build a strong encoder with developed Global-Local Spatial Perception (GLSP) blocks combining CNNs and RWKVs.
 We also propose a Cross-Channel Mix (CCM) module to improve skip connections with multi-scale feature fusion, achieving global channel information integration.
-Experiments on benchmark datasets, including Synapse, ACDC, BUSI, CVC-ClinicDB, CVC-ColonDB, Kvasir-SEG, ISIC 2017 and GLAS show that the RWKV-UNet achieves state-of-the-art performance on various types of medical image segmentation.
+Experiments on 11 benchmark datasets show that the RWKV-UNet achieves state-of-the-art performance on various types of medical image segmentation tasks.
 Additionally, smaller variants, RWKV-UNet-S and RWKV-UNet-T, balance accuracy and computational efficiency, making them suitable for broader clinical applications.
-1 Introduction
-Computer-aided medical image analysis plays a crucial role in modern healthcare.
-Automated image segmentation is a critical component of this process, as it divides complex images into distinct regions, allowing specific tissues, organs, or lesions to be more clearly visible, improving diagnostic accuracy, and facilitating the development of personalized treatment plans.
-The rise of deep learning [28] has greatly advanced image segmentation by automatically learning intricate patterns within images with deep neural networks, improving the precision and efficiency of the analysis.
-Current medical image segmentation approaches mainly rely on fully convolutional neural networks (FCNs) characterized by a U-shaped architecture.
-U-Net [46] is specifically designed for image segmentation tasks, particularly in the medical imaging domain.
-Its distinctive features include an encoder-decoder structure that enables effective downsampling and upsampling of images, along with skip connections.
-This design allows U-Net to capture both high-level semantic features and fine-grained details, resulting in improved segmentation accuracy.
-The architecture’s success has led to the development of numerous variants, such as Attention U-Net [40], U-Net++[61], U-Net3+ [22], which have further enhanced its performance and expanded its applicability.
-While CNNs excel at capturing local features through fixed-size kernels, they often struggle with long-range dependencies needed for understanding complex structures.
-Transformers, with their self-attention mechanism, capture global context and intricate patterns across entire images.
-Combining these architectures allows U-Net to retain fine detail, as seen in models like TransUNet [9], MT-UNet [53], and HiFormer[21].
-However, transformers have $O(N^{2})$ complexity in self-attention, which makes them computationally intensive for high-resolution scans.
-The trade-off between accuracy and efficiency is important in clinical scenarios, thus reducing computational complexity without significantly compromising accuracy is highly desirable.
-Recently, sequence modeling methods based on linear attention mechanisms, such as Mamba  [17] and RWKV  [41], have attracted significant attention and been applied to computer vision  [62, 16] and medical image segmentation tasks  [47, 38, 33].
-These models excel at capturing long-range dependencies with high accuracy, effectively preserving important contextual information across extensive image regions.
-In addition, their linear computational complexity greatly improves efficiency, making them suitable for large-scale or high-resolution images while reducing processing demands.
-Inspired by Vision-RWKV [47], we propose a novel RWKV-UNet, integrating convolutional layers with the spatial mix for medical image segmentation.
-The motivation behind this design is to take advantage of RWKV’s ability to capture long-range dependencies and combine it with the strengths of convolution localization, in order to achieve more precise and context-sensitive segmentation results.
-To evaluate our model’s effectiveness, we conducted experiments on benchmark datasets including Synapse  [27], ACDC  [5], BUSI  [1], CVC-ClinicDB  [4], CVC-ColonDB  [3], Kvasir-SEG  [25], ISIC 2017  [13] and GLAS  [49], where it achieves state-of-the-art (SOTA) performance, highlighting its potential as a robust tool for medical image analysis.
-In summary, our contributions are threefold.
-  • We propose a novel U-Net architecture based on a combination of RWKV and convolutional layers.
-An encoder based on the inverted residual RWKV (IR-RWKV) module is designed, effectively combining global and local information, with enhanced feature extraction capabilities through pre-training.
-  • A Cross-Channel Mix module is designed to improve the skip connections in U-Net, facilitating multi-scale information fusion and enhancing the model’s feature representation across different scales.
-  • Experiments conducted on datasets of different modalities show that the model achieves SOTA performance, demonstrating strong segmentation capabilities.
-We also propose relatively smaller models, RWKV-UNet-S and RWKV-UNet-T, which strike a balance between performance and computational efficiency, making them suitable for a wider range of scenarios.
-2 Related Work
-2.1 UNet Variants
-U-Net is a deep learning architecture designed for image segmentation, originally developed for biomedical images.
-Its symmetrical U-shaped structure consists of an encoder and a decoder, where the encoder progressively extracts features and reduces spatial dimensions, while the decoder gradually restores spatial information and generates high-resolution outputs.
-Skip connections are used to combine feature maps from the encoder layers with those from the decoder layers, ensuring the preservation of details.
-Variations in U-Net improve its performance through various means, including the use of deeper encoders such as ResNet  [59], the introduction of gating mechanisms to focus on important features  [40], improvements in skip connection formats  [61, 22, 64, 42] to recover more spatial details, as well as adaptations for 3D images [12, 39].
-Some studies have explored combining MLP/KAN with UNet to achieve long-range dependencies  [50, 35, 30].
+I Introduction
+Computer-aided medical image analysis is crucial in modern healthcare, where deep learning–driven automated segmentation precisely delineates anatomical structures and lesions, thereby enhancing visualization, improving diagnostic accuracy, and supporting personalized treatment planning.
+U-Net [1] is a seminal segmentation architecture featuring an encoder-decoder structure with skip connections, allowing it to capture both high-level semantic context and fine-grained details, which has inspired numerous improved variants.
+Medical images are typically high-resolution and exhibit complex anatomical structures with diverse lesion appearances, posing challenges for segmentation tasks that demand both fine local detail and coherent global context.
+Local features are essential for accurate boundary delineation and small lesion detection, whereas long-range dependencies ensure structural consistency and cross-slice contextual integration.
+CNN-based designs excel at local feature extraction through convolutional kernels but struggle to model long-range dependencies.
+Transformers-based [2, 3] UNet variants like [4, 5] address this via patch-based image processing and self-attention, improving segmentation accuracy, but the $O(N^{2})$ computational cost impacts efficiency, especially when processing high-resolution images.
+Figure 1: Comparative analysis of CNN-, Transformer-, Mamba-, RWKV-, and hybrid-based segmentation models, highlighting their respective strengths and weaknesses.
+Recent linear-attention models such as Mamba [6, 7] and RWKV [8, 9] have emerged as efficient alternatives, both achieving linear complexity while maintaining long-range modeling capability, and UNet variants based on these [10, 11, 12] have been developed for medical image segmentation.
+Mamba leverages state-space dynamics and selective scanning for efficient global information propagation, whereas RWKV introduces a gated recurrent mechanism that combines recurrence inductive bias with Transformer-style context aggregation.
+From a modeling perspective, RWKV’s gated accumulation enables more persistent and isotropic context propagation capability, yielding broader effective receptive fields and more stable long-range dependency modeling in medical image segmentation.
+Integrating CNNs with these global modeling mechanisms unifies precise local feature extraction and efficient global context understanding, addressing the inherent limitations of either CNNs or Transformers alone.
+Building on RWKV’s advantages, CNN–RWKV hybrids are expected to provide more balanced representations, offering a promising trade-off between segmentation accuracy and computational efficiency.
+Fig.
+1 compares the strengths and weaknesses of each model.
+Based on the advantages discussed above, we propose a novel RWKV-UNet for medical image segmentation, explicitly integrating convolutional layers with the spatial mixing capabilities of RWKV.
+This design leverages RWKV’s strength in capturing long-range dependencies and directionally consistent context propagation while retaining the precise local feature extraction of convolution, aiming to produce more accurate and context-aware segmentation results.
+We evaluate our model on 11 benchmark datasets, where it achieves state-of-the-art (SOTA) performance, demonstrating its effectiveness and potential as a robust tool for medical image analysis.
+In summary, our contributions are as follows:
+  • We propose an innovative RWKV-based Global-Local Spatial Perception (GLSP) module to build encoders for medical image segmentation.
+This approach efficiently integrates global and local information, and its feature extraction capabilities are augmented via pre-training.
+  • We design a Cross-Channel Mix (CCM) module to improve the skip connections in U-Net, which facilitates the fusion of multi-scale information and enhances the feature representation across different scales.
+  • Our RWKV-UNet demonstrates superiority and efficiency, achieving SOTA segmentation performance on 11 datasets of different imaging modalities.
+Additionally, the relatively compact models, namely RWKV-UNet-S and RWKV-UNet-T, strike a balance between performance and computational efficiency, making them adaptable to a wider range of application scenarios.
+II Related Work
+II-A Architectures for Medical Image Segmentation
+U-Net [1] is a deep learning architecture for biomedical image segmentation, featuring a U-shaped encoder-decoder structure with skip connections that preserve spatial details while extracting and restoring features.
+Variations based on U-Net improve its performance through various means, which include using developed encoders [13, 14], gating mechanisms to focus on important features [15], improving skip connection formats [16, 17, 18] to recover more spatial details, as well as adaptations for 3D images [19, 20].
+Some studies have explored combining MLP/KAN with UNet to capture long-range dependencies [21, 22, 23].
+nnU-Net [24] stands out as a self-configuring framework that automatically adapts preprocessing, architecture, and training settings to different medical datasets, serving as a robust baseline in many clinical segmentation challenges.
 These improvements allow U-Net to demonstrate higher accuracy and adaptability in various image segmentation tasks.
 Attention-based Improvements.
-Attention-based U-Net has several advantages, primarily its ability to capture global context information, which enhances the model’s understanding of long-range dependencies.
-Furthermore, the self-attention mechanism of Transformers allows the model to focus more effectively on key information.
-Variants of attention-based U-Net include pure Transformer models  [19, 23, 2, 8], which rely entirely on self-attention mechanisms for feature extraction and image segmentation.
-Another category consists of hybrid architectures that combine Transformers with CNNs  [9, 7, 53, 57, 52, 55, 21, 26], which leverage the local feature extraction capabilities of CNNs along with the global information processing strengths of Transformers to better capture image details and structural information.
-Recently, Mamba  [17]’s success in NLP and vision tasks  [62] has drawn significant attention to SSM, which can be considered a form of linear attention.
-SSM-based UNet variants  [47, 38, 54, 34, 33] capture long-range dependencies effectively and offer linear computational complexity, enhancing accuracy and efficiency in image segmentation tasks.
-2.2 Receptance Weighted Key Value (RWKV)
-Receptance Weighted Key Value (RWKV)  [41] was initially developed for natural language processing tasks.
-It combines the linear complexity of RNNs with the parallel processing advantages of Transformers, resulting in excellent performance, constant memory usage, and consistent inference generation speed.
-Building on this, Vision-RWKV (VRWKV)  [16] introduces RWKV to computer vision tasks, achieving success in image classification, object detection, and semantic segmentation.
-It efficiently handles sparse inputs and robust global processing and allows for high-resolution image processing without the need for window operations.
-Related efforts have also emerged in medical image analysis tasks.
-Restore-RWKV  [56] is an RWKV-based model for medical image restoration, enhancing the efficiency and effectiveness of medical image restoration through the introduction of the linear complexity Re-WKV attention mechanism and the omnidirectional token shift (Omni-Shift) layer.
-BSBP-RWKV [60, 60] achieves superior accuracy in medical image segmentation by combining the noise reduction advantages of Perona-Malik diffusion with the efficient RWKV structure.
-3 Methodology: RWKV-UNet
-3.1 Overall Architecture
+Attention-based methods have demonstrated strong performance in visual tasks due to their ability to capture global context.
+The self-attention mechanism of Transformers allows models to focus on the most informative regions, which is crucial in medical image segmentation for precise delineation of anatomical structures and lesions.
+Pure Transformer U-Nets [4, 5, 25, 26, 27] rely entirely on self-attention, enabling effective long-range dependency modeling but often at high computational cost.
+Hybrid CNN-Transformer architectures [28, 29, 30, 31, 32] combine CNNs’ local feature extraction with Transformers’ global modeling, preserving fine structural details while capturing complex tissue and lesion relationships.
+More recently, linear-attention sequence models such as Mamba [6, 7] and SSM-based U-Net variants [10, 11, 33, 34, 35] efficiently capture long-range dependencies with linear complexity, offering a favorable balance between segmentation accuracy and computational efficiency.
+And there are also attention-based attenpt for segmentation in medical videos [36].
+Despite the progress,existing parallel global-local fusion suffers from high computational cost and can interfere with fine feature representation, while some global modeling may be unnecessary in shallow layers.
+Additionally, existing approaches may struggle to capture long-range dependencies stably, as they lack recurrent inductive bias that support stable context propagation.
+II-B Receptance Weighted Key Value (RWKV)
+Receptance Weighted Key Value (RWKV) [8] models sequential dependencies via a weighted combination of past key-value pairs modulated by a learnable gate, avoiding the quadratic cost of traditional attention.
+Adapted to vision tasks as Vision-RWKV (VRWKV) [9], it efficiently captures long-range spatial dependencies in high-resolution images with linear complexity and reduced memory overhead, and has also been extended to broader visual domains [37, 38].
+RWKV for Medical Imaging.
+Restore-RWKV [39] applies linear Re-WKV attention with an Omni-Shift layer for efficient multi-directional information propagation in image restoration.
+BSBP-RWKV [40] integrates RWKV with Perona-Malik diffusion to achieve high-precision segmentation, preserving structural and pathological details.
+RWKVMatch [41] leverages Vision-RWKV-based global attention, cross-fusion mechanisms, and elastic transformation integration to handle complex deformations in medical image registration.
+Zigzag RWKV-in-RWKV (Zig-RiR) model [12] introduces a nested RWKV structure with Outer and Inner blocks and a Zigzag-WKV attention mechanism to capture both global and local features while preserving spatial continuity.
+These innovations demonstrate RWKV’s ability to capture long-range dependencies, retain fine anatomical features, and scale to high-resolution images.
+However, systematic exploration of hybrid architectures combining RWKV with local feature extractors remains limited, leaving opportunities to balance local detail, global context, and computational efficiency in medical image segmentation.
+III Methodology: RWKV-UNet
+III-A Overall Architecture
 Figure 2: Overall architecture of the proposed RWKV-UNet.
-(a) the encoder constructed by stacked IR blocks and stacked IR-RWKV blocks, (b) Cross-Channel Mix (CCM) Module, (c) the decoder constructed by point convolution layers and depthwise convolution layers,(d) the Inverted Residual (IR) block, (e) the Inverted Residual RWKV (IR-RWKV) block, (f) the architecture of the decoder block, containing a point-convolution layer and a $\times 9$ 9 DW-Conv layer, following a point operation layer and an upsampling operation.
-The overall architecture of the proposed RWKV-UNet is presented in Figure 1. RWKV-UNet consists of an encoder, a decoder and skip connections with a Cross-Channel Mix (CCM) Module.
-3.2 Effective Encoder in RWKV-UNet
-Inverted Residual RWKV (IR-RWKV) Block.
-We employ the spatial-mix module in Vision RWKV and Depth-Wise Convolution (DW-Conv) with a skip connection to combine global and local dependency, and trade-off model cost and accuracy.
+(a) the encoder with four stages constructed by stacked LP blocks and stacked GLSP blocks, (b) Cross-Channel Mix (CCM) Module for multi-scale fusion, (c) the decoder with four stages, (d) the Local Perception (LP) block, (e) the RWKV-based Global-Local Spatial Perception (GLSP) block, (f) the decoder block constructed by a point-convolution layer and a $9\times 9$ DW-Conv layer, with a convolution and an upsampling operation.
+The overall architecture of the proposed RWKV-UNet is presented in Fig.
+2.
+RWKV-UNet consists of an encoder with stacked LP and GLSP blocks, a decoder, and skip connections with a Cross-Channel Mix (CCM) Module.
+III-B Effective Encoder in RWKV-UNet
+Global-Local Spatial Perception (GLSP) Block
+We employ the RWKV-Spatial-Mix module and Depth-Wise Convolution (DW-Conv) with a skip connection to combine global and local dependency.
 The process of expanding and reducing dimensions enhances feature representation by capturing richer details while preventing information bottlenecks.
-Given a feature map $X\in\mathbb{R}^{C_{\mathrm{in}}\times H\times W}$, the processing flow through the IR-RWKV module is as follows:
-Normalize and project the input $X$ to an intermediate dimension $C_{mid}$ using a $\times 1$ 1 Convolution layer:
+Given a feature map $X\in\mathbb{R}^{C_{\mathrm{in}}\times H\times W}$, the processing flow through the GLSP module is as follows: Normalize and project the input $X$ to an intermediate dimension $C_{mid}$ using a $1\times 1$ Convolution layer:
 $I_{1}=\text{LayerNorm}(\text{Conv}_{1\times 1}(X)),$ (1)
 where $I_{1}\in\mathbb{R}^{C_{\mathrm{mid}}\times H\times W}$. $C_{\mathrm{mid}}$ should be greater than $C_{\mathrm{in}}$ to achieve dimension expansion.
-Divide the feature map into patches of size $\times 1$ 1. Then the Spatial Mix in VRWKV is applied:
+Divide the feature map into patches of size $1\times 1$. Then the Spatial Mix in VRWKV is applied:
 $I_{2}=\text{Unfolding}(I_{1}),$ (2)
@@
 where $I_{2}\in\mathbb{R}^{C_{\mathrm{mid}}\times N}$, $I_{3}\in\mathbb{R}^{C_{\mathrm{mid}}\times N}$.
-Convert the feature sequence back into a 2D feature map $I_{4}=\text{Folding}(I_{3})\in\mathbb{R}^{C_{\mathrm{mid}}\times H\times W}$, and use a $\times 5$ 5 DW-Conv layer for local feature aggregation:
+Convert the feature sequence back into a 2D feature map $I_{4}=\text{Folding}(I_{3})\in\mathbb{R}^{C_{\mathrm{mid}}\times H\times W}$, and use a $5\times 5$ DW-Conv layer for local feature aggregation:
 $I_{5}=\text{DW-Conv}(I_{4})+I_{4},$ (4)
@@
 where $F\in\mathbb{R}^{C_{\mathrm{out}}\times H^{\prime}\times W^{\prime}}$.
-Inverted Residual (IR) Block.
-An IR block contains a point convolution layer, a DW-Conv layer and a point convolution layer with local and global residual skips, removing the spatial mix as well as the unfolding and folding processes from the IR-RWKV.
-Effective Encoder with Stacked IRs and IR-RWKVs.
-We construct the RWKV-UNet encoder by stacking the IR and IR-RWKV blocks.
+Local Perception (LP) Block
+An LP block contains a point convolution layer, a DW-Conv layer and a point convolution layer with local and global residual skips, removing the spatial mix as well as the unfolding and folding processes from the GLSP.
+Effective Encoder with Stacked LPs and GLSPs
+We construct the RWKV-UNet encoder by stacking the LP and GLSP blocks.
 The encoder comprises a stem stage followed by four stages.
-The first blocks in stages \@slowromancapi@, \@slowromancapii@, \@slowromancapiii@ and \@slowromancapiv@ do not have residual connections and use a DW-Conv with a stride of 2 to achieve downsampling.
-After the first block, Stages \@slowromancapi@ and \@slowromancapii@ are composed of stacked IR blocks without Spatial Mix.
-In Stages \@slowromancapiii@ and \@slowromancapiv@, a series of IR-RWKV blocks follow the first block.
-Scale Up.
-As shown in Table 1, we implement three different sizes of encoders with different numbers of IR and IR-RWKV blocks in each stage (depths), different embedding dimensions and expansion ratios for $C_{\mathrm{mid}}$ : Encoder-Tiny (Enc-T), Encoder-Small (Enc-S), and Encoder-Base (Enc-B), allowing flexibility to scale up the model according to different resource constraints and application needs.
-Pre-training Manner.
+The first blocks in stages I, II, III and IV do not have residual connections and use a DW-Conv with a stride of 2 to achieve downsampling.
+After the first block, Stages I and II are composed of stacked LP blocks without Spatial Mix.
+In Stages III and IV, a series of GLSP blocks are stacked following the first block.
+Scale Up
+As shown in Table I, we implement three different sizes of encoders with different numbers of LP and GLSP blocks in each stage (depths), different embedding dimensions and expansion ratios for $C_{\mathrm{mid}}$ : Encoder-Tiny (Enc-T), Encoder-Small (Enc-S), and Encoder-Base (Enc-B), allowing flexibility to scale up the model according to different resource constraints and application needs.
+Pre-training Manner
 We pre-train the encoders of different sizes on ImageNet-1K [42] with a batch size of 1024 for 300 epochs using the AdamW [43] optimizer.
 These pre-trained weights of different sizes will be used in medical segmentation tasks, enabling improved feature extraction and faster convergence.
-Table 1: RWKV-UNet Encoder Configurations Across Stages
+TABLE I: RWKV-UNet Encoder Configurations Across Stages
 Stage Parameter Enc-T Enc-S Enc-B
 Stem Dimension 24 24 24
-Stage \@slowromancapi@ Depth 2 3 3
+Stage I Depth 2 3 3
 Dimension 32 32 48
@@
 Spatial Mix
-Stage \@slowromancapii@ Depth 2 3 3
+Stage II Depth 2 3 3
 Dimension 48 64 72
@@
 Spatial Mix
-Stage \@slowromancapiii@ Depth 4 6 6
+Stage III Depth 4 6 6
 Dimension 96 128 144
@@
 Spatial Mix
-Stage \@slowromancapiv@ Depth 2 3 3
+Stage IV Depth 2 3 3
 Dimension 160 192 240
@@
 Spatial Mix
-Parameters 2.6M 9.4M 16.7M
-FLOPs 1.9G 4.8G 9.0G
-Table 2: Comparison results of models with different attention mechanisms on Synapse dataset.
+Parameters 2.94M 9.42M 16.69M
+FLOPs 1.83G 4.73G 8.96G
+TABLE II: Comparison results of models with different attention mechanisms on Synapse dataset.
 Attention Type Average FLOPs
 HD95 $\downarrow$ DSC $\uparrow$
-No Attention 18.06 77.05 04.55G
-Multi-Head Self-attention [51] 16.95 77.60 14.89G
-Focused Linear Attention [18] 18.45 75.86 11.69G
-Bi-Mamba [62] 13.13 77.86 14.53G
-Spatial-Mix [47] 14.03 78.68 11.22G
+Multi-Head Self-Attention [2] 29.32 77.03 18.55G
+Focused Linear Attention [44] 27.66 76.89 11.18G
+Bi-Mamba [7] 29.10 78.13 14.53G
+RWKV Spatial-Mix (ours) [10] 25.01 78.14 11.11G
 Comparison with other attention types
-We conduct experiments by replacing the spatial mix in the encoder of RWKV-UNet with other attention mechanisms, such as self-attention  [51], focused linear attention  [18] and Bi-Mamba used in Vision Mamba  [62], as well as by removing this module entirely.
-The total training epochs are 60.
-The results shown in Table 2 indicate that the model with Spatial-Mix achieves the best DSC with the lowest computational cost compared to other attention mechanisms, showcasing its efficiency and effectiveness.
-And all kinds of attention mechanisms can lead to improvement compared to pure CNN structure except focused linear attention.
-3.3 Decoder Design
-We design our CNN-based decoder block with a point-convolution layer and a $\times 9$ 9 DW-Conv layer, following a point operation layer and an upsampling operation.
+We conduct experiments by replacing the spatial mix in the encoder of RWKV-UNet with other attention mechanisms, such as self-attention [2], focused linear attention [44], and Bi-Mamba used in Vision Mamba [7], as well as by removing this module entirely.
+The total training epochs are 100.
+The results shown in Table II indicate that the model with Spatial-Mix achieves the best DSC with the lowest computational cost.
+Comparison visualizations of effective receptive fields of the last layer output using different attention mechanisms in the GLSP module are shown in Fig.
+3.
+(a) No Attention
+(b) MHSA
+(c) Bi-Mamba
+(d) Spatial Mix
+Figure 3: Comparison visualization of effective receptive fields of the last layer output using different attention mechanisms in the GLSP module.
+III-C Large Kernel-based Decoder Design
+We design our CNN-based decoder block with a point-convolution layer and a $9\times 9$ DW-Conv layer, following a point operation layer and an upsampling operation.
 Consider the input feature map $X\in\mathbb{R}^{C_{\text{in }}\times H\times W}$; in depthwise convolution, the number of input channels is $C_{\mathrm{in}}$, and the second point convolution maps the $C_{\text{in }}$ channels to $C_{\text{out }}$.
-The computational complexity of a standard convolution is:
-$H\times W\times C_{\text{in }}\times\left(C_{\text{out }}\times K^{2}\right).$ (6)
-The computational complexity of pointwise convolution + depthwise convolution + pointwise convolution is:
-$H\times W\times C_{\text{in }}\times\left(C_{\text{in }}+K^{2}+C_{\text{out }}\right),$ (7)
-where $K$ is the kernel size.
-The latter method is more computationally efficient for the use of larger kernels, as the complexity is proportional to $C_{\text{in }}\times K^{2}$ rather than $C_{\text{in }}\times C_{\text{out }}\times K^{2}$.
-3.4 Cross-Channel Mix for Multi-scale Fusion
+III-D Cross-Channel Mix for Multi-scale Fusion
 Inspired by the Channel Mix in VRWKV, we propose Cross-Channel Mix (CCM), a module that can effectively extract channel information across multi-scale encoder features.
 By capturing the rich global context along channels, CCM further enhances the extraction of long-range contextual information.
-Define output feature map of Stage \@slowromancapi@, Stage \@slowromancapii@ and Stage \@slowromancapiii@ of the designed encoder are $F_{1},F_{2}$, and $F_{3}$, which has different Spatial dimensions and channel counts: $F_{1}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}},\quad F_{2}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}},\quad F_{3}\in\mathbb{R}^{C_{3}\times H_{3}\times W_{3}}$.
-($H_{1}$ >$H_{2}$>$H_{3}$, $W_{1}$ >$W_{2}$ >$W_{3}$ and $C_{3}$ >$C_{2}$ >$C_{1}$).
+Define output feature map of Stage I, Stage II and Stage III of the designed encoder are $F_{1},F_{2}$, and $F_{3}$, which has different Spatial dimensions and channel counts: $F_{1}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}},\quad F_{2}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}},\quad F_{3}\in\mathbb{R}^{C_{3}\times H_{3}\times W_{3}}$.
+($H_{1}$ $H_{2}$$H_{3}$, $W_{1}$ $W_{2}$ $W_{3}$ and $C_{3}$ $C_{2}$ $C_{1}$).
 We reshape and map smaller features to match the largest feature’s size and dimension:
 $\begin{gathered}\tilde{F}_{i}=\operatorname{Conv}_{C_{i}}\left(\operatorname{Upsample}\left(F_{i}\right)\right),\quad i=2,3\\ \tilde{F}_{1}=\operatorname{Conv}_{C_{1}}\left(F_{1}\right),\end{gathered}$ (6)
-where $\tilde{F}_{1}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$, $\tilde{F}_{2}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$, and $F_{3}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$.
+where $\tilde{F}_{i}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}},\quad i=1,2,3$
 We concatenate the adjusted feature maps along the channel dimension to produce a combined feature map:
@@
 where $F_{1}^{\prime}\in\mathbb{R}^{C_{1}\times H_{1}\times W_{1}}$ , $F_{2}^{\prime}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$ and $F_{3}^{\prime}\in\mathbb{R}^{C_{3}\times H_{3}\times W_{3}}$.
-$F_{1}^{\prime}$,$F_{2}^{\prime}$ and $F_{3}^{\prime}$ will be concatenated with the output of the decoder Stage \@slowromancapiii@, Stage \@slowromancapii@ and Stage \@slowromancapi@.
-4 Experiments
-Figure 2: Average DSC vs. #Params for different methods on Synapse multi-organ segmentation dataset.
-Figure 3: Average DSC vs. #FLOPs for different methods on Synapse multi-organ segmentation dataset.
-[…]
+$F_{1}^{\prime}$$F_{2}^{\prime}$ and $F_{3}^{\prime}$ will be concatenated with the output of the decoder Stage III, Stage II and I.
+IV Experiments
+TABLE III: Comparison results on Synapse dataset.
+The evaluation metrics are HD95 ($m$) and DSC in (%).
+DSC are reported for individual organs.
+$\uparrow\downarrow$ denotes the higher (lower) the better.
+- means missing data from the source.
+Bold and underline represent the best and the second best results.
+denotes that the experiment is conducted by us.
 Methods Average Aotra Gallbladder Kidney (Left) Kidney (Right) Liver Pancreas Spleen Stomach
@@
 R50 U-Net [45, 1] 36.87 74.68 87.74 63.66 80.60 78.19 93.74 56.90 85.87 74.16
-R50 Att-Unet  [20, 40] 36.97 75.57 55.92 63.91 79.20 72.71 93.56 49.37 87.19 74.95
 ViT [3] + CUP [28] 36.11 67.86 70.19 45.10 74.70 67.40 91.32 42.00 81.75 70.44
-R50-ViT  [20, 15] + CUP [9] 32.87 71.29 73.73 55.13 75.80 72.20 91.51 45.99 81.99 73.95
 TransUNet [28] 31.69 77.48 87.23 63.16 81.87 77.02 94.08 55.86 85.08 75.62
@@
 MISSFormer [5] 18.20 81.96 86.99 68.65 85.21 82.00 94.41 65.67 91.92 80.81
-CASTformer [57] 22.73 82.55 89.05 67.48 86.05 82.17 95.61 67.49 91.00 81.55
 TransDeepLab [25] 21.25 80.16 86.04 69.16 84.08 79.88 93.53 61.19 89.00 78.40
 LeVit-Unet-384 [48] 16.84 78.53 87.33 62.23 84.61 80.25 93.11 59.07 88.86 72.76
-CTC-Net  [58] - 78.41 86.46 63.53 83.71 80.79 93.78 59.73 86.87 72.39
 MS-UNet [26] 18.97 80.44 85.80 69.40 85.86 81.66 94.24 57.66 90.53 78.33
@@
 TransCASCADE [49] 17.34 82.68 86.63 68.48 87.66 84.56 94.43 65.33 90.79 83.52
-AiA-UNet  [43] - 82.88 87.77 74.33 84.16 82.39 94.98 65.65 92.77 80.99
 MetaSeg-B [32] - 82.78 - - - - - - - -
 VM-UNet [10] 19.21 81.08 86.40 69.41 86.16 82.76 94.17 58.80 89.51 81.40
+Hc-Mamba [50] 26.34 79.58 89.93 67.65 84.57 78.27 95.38 52.08 89.49 79.84
+Mamba-UNet [33] 17.32 73.36 81.65 55.39 81.86 72.76 92.35 45.28 88.22 69.35
 PVT-EMCAD-B2 [51] 15.68 83.63 88.14 68.87 88.08 84.10 95.26 68.51 92.17 83.92
-SelfReg-UNet  [63, 46] - 80.34 88.74 71.78 85.32 80.71 93.80 62.22 84.78 75.39
 SelfReg-SwinUnet [52, 46] - 80.54 86.07 69.65 85.12 82.58 94.18 61.08 87.42 78.22
-RWKV-UNet-T (ours) 17.83 81.58 89.28 64.18 86.03 83.76 95.09 63.50 90.36 80.44
-RWKV-UNet-S (ours) 21.16 82.20 89.36 67.60 84.11 79.26 95.60 68.86 91.58 81.21
-RWKV-UNet (ours) 15.70 84.02 89.53 68.94 87.63 84.07 95.57 69.38 90.95 86.09
-Table 4: Comparison results on ACDC dataset. The evaluation metric is DSC in (%).
+Zig-RiR (2D) [12] ($\times256$) 28.51 74.61 82.73 64.19 79.13 68.25 93.71 50.14 88.46 70.29
+RWKV-UNet-T (ours) 11.89 81.87 88.10 68.42 88.88 82.61 95.15 60.52 90.23 81.09
+RWKV-UNet-S (ours) 11.01 82.93 88.30 65.92 88.47 84.11 95.45 66.51 91.51 83.17
+RWKV-UNet (ours) 08.85 84.29 88.71 70.39 89.58 84.36 95.70 66.58 93.54 85.49
+TABLE IV: Comparison results on ACDC dataset. The evaluation metric is DSC in (%).
 Methods Average RV Myo LV
@@
 R50-ViT [45, 3] + CUP [28] 87.57 86.07 81.88 94.75
-R50 Att-UNet  [20, 40] 86.75 87.58 79.20 93.47
 TransUNet [28] 89.71 88.86 84.54 95.73
@@
 MISSFormer [5] 90.86 89.55 88.04 94.99
-CTC-Net  [58] 90.77 90.09 85.52 96.72
 LeViT-UNet-384s [48] 90.32 89.55 87.64 93.76
 PVT-CASCADE [49] 90.45 87.20 88.96 95.19
+VM-UNet [10] 91.47 89.93 89.04 95.44
 TransCASCADE [49] 91.63 89.14 90.25 95.50
 PVT-EMCAD-B2 [51] 92.12 90.65 89.68 96.02
-SelfReg-UNet  [63, 46] 91.43 88.92 89.49 95.88
 SelfReg-SwinUnet [52, 46] 91.49 89.49 89.27 95.70
-RWKV-UNet-T (ours) 90.92 89.73 87.18 95.85
-RWKV-UNet-S (ours) 91.64 89.77 88.48 96.65
-RWKV-UNet (ours) 92.17 90.86 88.72 96.92
-Table 5: Comparison results on BUSI, CVC-ClinicDB, CVC-ColonDB, Kvasir-SEG, ISIC 2017 and GLAS datasets.
+Zig-RiR (2D) [12] ($\times256$) 87.42 86.41 81.87 93.98
+RWKV-UNet-T (ours) 91.90 90.63 88.21 96.85
+RWKV-UNet-S (ours) 91.19 89.49 87.87 96.22
+RWKV-UNet (ours) 92.29 91.26 88.78 96.83
+TABLE V: Comparison results on GOALS and FUGC 2025 dataset. The evaluation metric is DSC in (%). All experiments for baselines are conducted by us.
+Methods GOALS FUGC 2025
+U-Net [1] 90.99 51.96
+TransUNet [28] 91.05 79.87
+Rolling-UNet-M [22] 90.84 63.39
+PVT-EMCAD-B2 [51] 90.91 57.39
+PVTFormer [53] 90.98 78.41
+UKAN [23] 90.59 72.19
+VM-UNet [10] 91.10 72.94
+H-vmunet [54] 84.18 65.99
+Zig-RiR (2D) [12] 83.48 62.41
+RWKV-UNet-T (ours) 91.34 78.67
+RWKV-UNet-S (ours) 91.13 79.43
+RWKV-UNet (ours) 91.75 81.17
+TABLE VI: Comparison results on BUSI, CVC-ClinicDB, CVC-ColonDB, Kvasir-SEG, ISIC 2017, GLAS and PSVFM datasets.
 The evaluation metric are DSC in (%).
-Results with error bars are averaged over three runs with three different splits.
-Methods BUSI CVC-ClinicDB CVC-ColonDB Kvasir-SEG ISIC 2017 GLAS Params. FLOPs
-U-Net [46] 74.90$\pm$1.15 89.22$\pm$4.07 83.43$\pm$5.24 84.94$\pm$2.60 82.94 88.56$\pm$0.63 07.8M 12.15G
-Att-UNet [40] 78.14$\pm$1.25 90.86$\pm$4.09 84.75$\pm$6.55 86.21$\pm$2.15 83.96 89.14$\pm$0.54 08.7M 16.79G
-UNet++ [61] 75.74$\pm$1.70 90.07$\pm$5.09 84.75$\pm$3.99 85.86$\pm$1.52 82.27 88.93$\pm$0.67 09.2M 34.77G
-TransUNet [9] 79.35$\pm$2.45 93.91$\pm$2.28 90.72$\pm$3.92 91.34$\pm$0.01 86.09 91.58$\pm$0.15 00.1G 38.54G
-FCBFormer [48] 80.66$\pm$0.28 91.15$\pm$3.50 85.23$\pm$5.21 91.09$\pm$0.96 84.56 91.13$\pm$0.25 53.0M 41.60G
-PVT-CASCADE  [44] 78.30$\pm$3.17 93.45$\pm$1.67 90.97$\pm$2.11 91.40$\pm$1.29 84.77 91.85$\pm$0.39 35.3M 08.40G
-Rolling-UNet-M  [31] 77.49$\pm$2.42 91.91$\pm$2.56 85.20$\pm$5.85 87.41$\pm$1.78 84.42 88.12$\pm$0.66 07.1M 08.31G
-PVT-EMCAD-B2  [45] 79.22$\pm$1.07 93.39$\pm$1.64 91.27$\pm$1.84 90.37$\pm$1.10 84.05 91.77$\pm$0.09 26.8M 05.85G
-U-KAN  [30] 76.95$\pm$1.67 89.88$\pm$3.21 84.66$\pm$5.33 85.78$\pm$1.68 83.46 87.42$\pm$0.28 25.4M 10.59G
-RWKV-UNet-T (ours) 79.47$\pm$1.35 95.05$\pm$0.67 89.56$\pm$5.26 91.26$\pm$2.13 85.59 91.96$\pm$0.31 03.3M 03.78G
-RWKV-UNet-S (ours) 80.92$\pm$1.74 95.58$\pm$0.69 91.30$\pm$3.21 91.78$\pm$1.24 86.38 92.11$\pm$0.16 09.9M 07.75G
-RWKV-UNet (ours) 81.92$\pm$0.82 95.26$\pm$0.70 92.27$\pm$2.48 91.26$\pm$1.27 85.32 92.35$\pm$0.07 17.4M 14.64G
-4.1 Experimental Setup
-Datasets and Metrics.
-Experiments on conducted on Synapse [27] for multi-organ Segmentation in CT images, ACDC [5] for cardiac segmentation in MRI images, BUSI [1] for breast tumor segmentation in ultrasound images, CVC-ClinicDB [4], CVC-ColonDB [3], Kvasir-SEG [25] for poly segmentation, ISIC 2017 [13] for skin lesion segmentation and GLAS  [49] for gland segmentation.
+All experiments for baselines are conducted by us.
+Results with error bars are averaged over three runs with different splits.
+Methods BUSI CVC-ClinicDB CVC-ColonDB Kvasir-SEG ISIC 2017 GLAS PSVFM Params. FLOPs
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 4: Performance of different methods on the Synapse multi-organ segmentation dataset.
+The average DSC (%) is plotted against FLOPs (G).
+The size of each circle represents the model’s parameter count.
+RWKV-UNet achieves SOTA performance with balanced computation cost, while RWKV-UNet-S and RWKV-UNet-T also achieve remarkable results.
+IV-A Experimental Setup
+Datasets and Metrics
+Experiments on conducted on Synapse [56] for multi-organ Segmentation in CT images, ACDC [57] for cardiac segmentation in MRI images, GOALS [58] for OCT layer segmentation, BUSI [59] for breast tumor segmentation in ultrasound images, CVC-ClinicDB [60], CVC-ColonDB [61], Kvasir-SEG [62] for poly segmentation in endoscopy images, ISIC 2017 [63] for skin lesion segmentation in dermoscopic images, GLAS [64] for gland segmentation in microscopy images, PSVFM [65] for placental vessel segmentation in fetoscopy images and FUGC [66] for semi-supervised cervix segmentation in ultrasound images.
 The average Dice Similarity Coefficient (DSC) and average 95% Hausdorff Distance (HD95) are used as evaluation metrics.
-Specific details for datasets can be seen in Appendix.
-Comparison Methods.
-We use U-Net  [46], R50 U-Net  [20, 46], R50 Att-Unet  [20, 40], ViT  [15] + CUP  [9], R50-ViT  [20, 15] + CUP  [9], TransUNet  [9], SwinUNet  [6], TransClaw U-Net  [7], MT-UNet  [53], UCTransNet  [52], MISSFormer  [23], CASTformer  [57], TransDeepLab  [2], CTC–Net  [58], MS-UNet  [8], LeVit-Unet-384  [55], HiFormer-L  [21], PVT-CASCADE  [44], TransCASCADE  [44], AiA-UNet  [43], MetaSeg-B  [26], VM-UNet  [47],PVT-EMCAD-B2  [45], SelfReg-UNet  [63, 46], SelfReg-SwinUnet  [63, 6], FCBFormer [48], Rolling-UNet-M  [31], U-KAN  [30]as baselines.
-Implementation Details.
-All experiments are all performed on PG500-216(V-100) with 32 GB memory.
-The resolution of the input images is resized to 224$\times$224 for Synapse and ACDC, and 256$\times$256 for binary segmentation tasks.
-The total training epochs for Synapse are 30, for ACDC are 150, and for binary segmentation are 300.
-The batch size is 24 for Synapse and ACDC, and 8 for binary segmentation tasks.
-The initial learning rate is 1e-3 for Synapse, 5e-4 for ACDC, and 1e-4 for binary segmentation experiments.
-The minimum learning rates for Synapse and ACDC are 0 and for binary segmentation are 1e-5.
-The AdamW  [36] optimizer and CosineAnnealingLR  [37] scheduler are used.
-Test-time augmentation is used for Synapse and ACDC.
-The baseline results on binary segmentation tasks are run by ourselves.
-The loss function used is a mixed loss combining cross entropy (CE) loss and dice loss  [39]:
+  • Synapse uses 18 training and 12 validation CT scans, while ACDC uses 70 training, 10 validation, and 20 testing MRI scans, following settings in [28].
+  • GOALS uses 100 training, 100 validation and 100 test images following the original setting in the challenge.
+  • BUSI, CVC-ClinicDB, CVC-ColonDB, and Kvasir-SEG are split into training, validation, and testing sets with an 8:1:1 ratio, repeated with three different random seeds.
+  • ISIC 2017 uses 2000 training, 150 validation, and 600 testing images as per the original challenge setup.
+  • GlaS uses 85 training and 80 testing whole-slide images, with the training set further split into training and validation with a 9:1 ratio across three random seeds.
+  • PSVFM uses a video-based split: Videos 1–3 for training, Video 4 for validation, and Videos 5–6 for testing, reflecting real-case variability in fetoscopic surgeries.
+  • Building on [67] and [67], FUGC uses 50 labeled and 450 unlabeled images for training, 90 for validation, and 300 for testing.
+Implementation Details
+[…]
 $\mathcal{L}=\alpha CE(\hat{y},y)+\beta Dice(\hat{y},y),$ (10)
-where $\alpha$ and $\beta$ are 0.5 and 0.5 for Synapse, 0.3 and 0.7 for ACDC, 0.5 and 1 for binary segmentation.
+where $\alpha$ and $\beta$ are 0.5 and 0.5 for Synapse, 0.2 and 0.8 for ACDC, 0.5 and 1 for GOALS and binary segmentation.
 Experiments for PVT-CASCADE [49] and PVT-EMCAD-B2 [51] use deep supervision strategy [69].
+For semi-supervised tasks, the supervised loss on labeled data remains $\mathcal{L}_{\text{sup}}=0.5CE(\hat{y},y)+Dice(\hat{y},y)$.
+In addition, a consistency loss is applied on unlabeled data, computed as the mean squared error (MSE) between predictions under different augmentations after inverting the transformations.
+Learning rate for TransUNet and PVTFormer are $1e^{-4}$ on FUGC 2025 for better results.
+IV-B Comparison with State-of-the-arts
+Abdomen Multi-organ Segmentation
+Table III shows that RWKV-UNet excels in abdomen organ segmentation on the Synapse dataset, achieving the highest average DSC of 84.29% , and surpassing all SOTA CNN-, transformer-, and Mamba-based methods.
+It also achieves 8.85 in HD95, which is significantly better than other models, demonstrating a strong ability to localize organ boundaries accurately.
+This can be attributed to the model’s ability to capture both long-range dependencies and local features.
+Smaller models, RWKV-UNet-T and RWKV-UNet-S, also achieve remarkable results.
+The DSCs with parameters and FLOPs of each model can be seen in Fig.
+4 and visualization results can be seen in Fig.
+5.
+Cardiac Segmentation
+Table IV shows that our model outperforms all SOTA models in the ACDC MRI cardiac segmentation task, achieving the best results in RV and the second-best in LV categories.
+The smaller models, RWKV-UNet-T and RWKV-UNet-S, also perform well.
+OCT Layer Segmentation
+Table V shows that on the small GOALS dataset, performance differences are minor and some carefully designed models even lag behind the classic U-Net, while our RWKV-UNet still achieves the highest average DSC, demonstrating its robustness and superiority.
+These results also reflect the model’s ability to handle small datasets.
+Binary Segmentation
+Table VI shows that the RWKV-UNet series shows remarkable performance in the tasks of breast lesion, poly, skin lesion, grand, and vessel segmentation, with much smaller parameters and computation load than TransUNet and FCBFormer.
+Semi-supervised Cervix Segmentation
+Table V also shows that the RWKV-UNet achieves SOTA performance on the FUGC 2025 dataset, further indicating strong data efficiency and generalization ability, as it can fully leverage limited labeled data and mine useful information from unlabeled data to maintain high segmentation accuracy.
 Figure 5: A qualitative comparison with previous SOTA methods on the Synapse dataset.
 The visual results demonstrate that our method achieves more accurate segmentation, especially in difficult tasks like pancreas segmentation.
-4.2 Comparison with State-of-the-arts
-Multi-organ Segmentation.
-Table 3 shows that RWKV-UNet excels in abdomen organ segmentation on the Synapse multi-organ dataset, achieving the highest average DICE score of 84.02% and surpassing all SOTA CNN- and transformer-based methods.
-It also achieves 15.70 in HD95, only slightly lower than PVT-EMCAD-B2, demonstrating a strong ability to localize the organ boundary.
-Its performance on the most challenging task, pancreas segmentation, is the best, indicating its effectiveness in handling sample imbalance issues.
-Smaller models, RWKV-UNet-T and RWKV-UNet-S, also achieve remarkable results.
-Cardiac Segmentation.
-Table 4 shows that our model outperforms all SOTA models in the ACDC MRI cardiac segmentation task, achieving the best results in both RV and LV categories.
-The smaller models, RWKV-UNet-T and RWKV-UNet-S, also perform well.
-Binary Segmentation.
-Table 5 shows that RWKV-UNet series show remarkable performance in the tasks of breast lesion, poly, skin lesion and grand segmentation, with much smaller parameters and computation load than TransUNet and FCBFormer.
-Meanwhile, it is observed that in some datasets the smaller model RWKV-UNet-S perform better.
-4.3 Ablation and Analysis
-Effect of Pre-training for Encoder.
-Whether to use pre-trained weights of RWKV-UNet encoder results in Table 6.
-Experimental results indicate that pre-training is essential because it significantly improves feature extraction by enabling the encoder to capture intricate patterns and hierarchical information.
-Notably, without pre-training, HD95 is higher.
-But DSC is a more important metric, providing a reliable assessment of segmentation quality directly, while HD95 can be overly sensitive to outliers.
-Table 6: Comparison results on Synapse dataset of whether to use pre-trained weights on ImageNet-1k or not.
+IV-C Ablation Study and Additional Analysis
+Comparison with Different Encoders
+We replace the encoder with pre-trained weights while keeping the rest of the network architecture unchanged (dimensions vary according to the encoder).
+Table VII shows that our pretrained encoder achieves the best DSC and HD95 on the Synapse dataset, with significantly lower FLOPs.
+TABLE VII: Comparison results on Synapse dataset of different pretrained encoders
+Encoder Average FLOPs
+HD95 $\downarrow$ DSC $\uparrow$
+ResNet 50 [45] 9.99 82.27 62.44
+PVT-B3-v2 [70] 11.47 81.23 22.74
+PVT-B5-v2 [70] 10.53 80.14 27.38
+ConvNext-small [71] 12.79 76.34 42.47
+ConvNext-base [71] 10.61 81.42 74.86
+MaxViT-small [72] 10.57 83.71 18.54
+MaxViT-base [72] 11.09 82.89 29.86
+RWKV-UNet Enc-B (ours) 08.85 84.29 11.11
+Effect of Pre-training for Encoder
+Experimental results in Table VIII.
+indicate that pre-training is essential because it significantly improves feature extraction by enabling the encoder to capture intricate patterns and hierarchical information.
+And the inductive biases of RWKV are also not as strong as those of CNN, thus they also rely on pre-training similar to transformers.
+TABLE VIII: Comparison results on Synapse dataset of whether to use pre-trained weights on ImageNet-1k or not.
 Pre-training Training Epochs Average
 HD95 $\downarrow$ DSC $\uparrow$
-w/o 30 13.94 76.49
-w/o 60 14.03 78.68
-w/ 30 15.70 84.02
-Table 7: Comparison results of different decoder designs for RWKV-UNet on Synapse dataset.
-$c\rightarrow c\rightarrow d$ denotes $\operatorname{Conv}_{1\times 1}\rightarrow\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv and $c\rightarrow d\rightarrow c$ denotes $\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv $\rightarrow\operatorname{Conv}_{1\times 1}.$
-Decoder Kernel Size Average Flops
+w/o 30 36.23 71.66
+w/o 100 25.01 78.14
+w/ 30 08.85 84.29
+TABLE IX: Comparison results of different decoder designs for RWKV-UNet on Synapse dataset.
+Kernel Size Synapse ACDC FLOPs
+HD95 $\downarrow$ DSC $\uparrow$ DSC $\uparrow$
+3 10.35 84.83 92.07 10.97G
+5 08.69 84.67 91.95 11.00G
+7 09.97 83.73 91.61 11.05G
+9 08.85 84.29 92.29 11.11G
+11 11.25 82.67 92.04 11.19G
+Attention in Shallow Stages
+We evaluate the effect of retaining attention in the first two layers of RWKV-UNet-S by replacing them with spatial-mix attention and comparing against the original CNN-based shallow layers.
+All models are trained for 100 epochs from scratch, without pretraining.
+Table X shows negligible improvement from shallow attention on Synapse, while introducing higher computational cost, indicating that CNNs are sufficient for capturing low-level features in the shallow layers.
+TABLE X: Comparison of results on the Synapse dataset: effect of attention in shallow stages of RWKV-UNet-S
+Shallow Attention Average FLOPs
 HD95 $\downarrow$ DSC $\uparrow$
-Conv + upsample 3 24.95 81.77 11.38G
-$c\rightarrow c\rightarrow d$
-+ upsample
-3 23.44 82.10 11.06G
-5 25.53 81.84 11.07G
-7 20.78 81.57 11.08G
-9 18.07 83.76 11.10G
-$c\rightarrow d\rightarrow c$
-+ upsample
-3 21.68 83.33 11.08G
-5 22.26 83.74 11.11G
-7 22.90 82.54 11.16G
-9 15.70 84.02 11.22G
-Decoder Design.
-In Tabel 7, we examine various sizes of convolution kernels, finding that the use of large kernels, such as $\times 9$ 9, in the decoder significantly enhances the segmentation performance.
-This improvement is likely due to larger kernels capturing broader contextual information and preserving more spatial details, which enhances segmentation accuracy in medical images.
-We also compare different ways to construct the decoder block and the results show that $\operatorname{Conv}_{1\times 1}\rightarrow\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv performs better than $\operatorname{Conv}_{1\times 1}\rightarrow$ DW-Conv $\rightarrow\operatorname{Conv}_{1\times 1}$ and the common convolution layer, showing that DW-Conv operating on larger dimensions will help the decoding process.
-Channel Mix.
+w/ 29.43 76.88 6.91G
+w/o (ours) 28.92 76.41 5.86G
+Channel Mix in GLSP Blocks
 In VRWKV [9], a Channel Mix module is used after the Spatial Mix.
-In RWKV-UNet’s encoder, IR-RWKV blocks are designed without using Channel Mix as the pointwise convolution in the output layer performs both channel mix and feed-forward operations.
-We compare the results on the Synapse dataset with and without applying channel mix after spatial mix.
-All experiments are performed without pre-training, with a total of 60 epochs.
-The results are shown in Table 8, demonstrating that the channel mix is not essential for maintaining performance.
-Table 8: Comparison of results on the Synapse dataset: effect of applying Channel Mix after Spatial Mix in the IR-RWKV blocks
+In contrast, our RWKV-UNet’s encoder designs GLSP blocks without Channel Mix, as the pointwise convolution in the output layer performs both channel mixing and feed-forward operations.
+We compare results on the Synapse dataset (without pre-training, 100 epochs) .
+Table XI demonstrates that Channel Mix is not essential for maintaining performance.
+TABLE XI: Comparison of results on the Synapse dataset: effects of Channel Mix after Spatial Mix in GLSP blocks
 Channel Mix Average FLOPs
 HD95 $\downarrow$ DSC $\uparrow$
-w/ 15.29 78.76 19.49G
-w/o (ours) 14.03 78.68 11.22G
-Skip Connections.
-Comparison results of different skip connection designs for RWKV-UNet on the Synapse dataset are shown in Table 9.
+w/ 35.19 74.41 19.43G
+w/o (ours) 25.01 78.14 11.11G
+Skip Connections
+Comparison results of different skip connection designs for RWKV-UNet on the Synapse dataset are shown in Table XII.
 Experiments analyze variations in the number of skip connections and the inclusion of the CCM block.
@@
 Furthermore, the results show that the CCM module can significantly improve the model segmentation results.
-Even if only one or two inputs from the encoder(when only one input, the CCM block degrades to a ”channel mix bridge”), the CCM block can still lead to increments in DSC.
+With only two skips, the model can still achieve 84.28 DSC with CCM.
 Admittedly, the CCM module increases the computational load to some extent due to the aggregation of global information.
-Table 9: Comparison results on different skip connections design of RWKV-UNet on Synapse dataset. Experiments are conducted on the number of skips and whether or not to use the CCM block.
+TABLE XII: Comparison results on different skip connections designs of RWKV-UNet on the Synapse dataset. Experiments are conducted on the number of skips and whether or not to use the CCM block.
 Skips CCM Block Average FLOPs
 HD95 $\downarrow$ DSC $\uparrow$
-0 w/o 23.39 74.19 09.25 G
-1 w/o 17.17 79.76 09.41 G
-1 w/ 19.56 79.98 09.63 G
-2 w/o 16.08 80.90 09.49 G
-2 w/ 19.97 81.11 10.24 G
-3 w/o 22.72 81.14 09.56 G
-3 w/ 15.70 84.02 11.22 G
-5 Conclusion
-In this study, we introduce RWKV-UNet, a novel architecture that integrates the RWKV architecture with U-Net.
-By combining the strengths of convolutional networks for local feature extraction and RWKV’s ability to model global context, our model significantly improves medical image segmentation accuracy.
-The proposed enhancements, such as the IR-RWKV module and the CCM module, contribute to a more precise feature representation and information fusion across different scales.
-Experimental results on benchmark datasets demonstrate that RWKV-UNet surpasses all SOTA CNN- and transformer-based methods.
-Its variants (RWKV-UNet-S and RWKV-UNet-T) offer a practical balance between performance and computational efficiency.
-Our approach has a strong potential to advance medical image analysis, particularly in clinical settings where both accuracy and efficiency are paramount.
-Limitations and Future Work.
-RWKV-UNet is a powerful 2D medical image segmentation model that effectively combines RWKV with convolution, but it is currently not applicable to 3D imaging.
-In the future, we plan to extend the model to 3D for handling volume data and to explore the potential of RWKV-based foundational models for medical imaging.
-We also aim to develop ultra-lightweight RWKV-based models tailored for ’bedside healthcare’ applications, preserving segmentation accuracy while enhancing adaptability and speed further.
-Appendix
-The supplementary material gives more specific details of our experimental settings, and presents more comprehensive analysis and results of our RWKV-UNet to facilitate the comparison of subsequent methods:
-  • Sec. A provides introduction to datasets and corresponding splitting settings used in this paper.
-  • Sec. B.1 provides comparison results on Synapse dataset to explore the effect of applying DW-Conv after Spatial Mix in IR-RWKV blocks.
-  • Sec. B.2 provides comparison results of different methods on Synapse dataset with larger inputs.
-Appendix A Introduction to datasets
-1) The Synapse multi-organ segmentation dataset  [27]: 30 abdominal CT scans from the MICCAI 2015 Multi-Atlas Abdomen Labeling Challenge are used.
-The dataset is split into 18 training cases and 12 cases for validation following the setting in  [9].
-2) The Automated Cardiac Diagnosis Challenge (ACDC)  [5]: This dataset gathezrs examinations obtained by MRI scanners.
-Each patient scan is manually annotated with ground truth data for the left ventricle (LV), right ventricle (RV), and myocardium (MYO).
-The dataset is split into 70 training cases, along with 10 cases for validation and 20 for testing following the setting in  [9].
-3) Breast Ultrasound Images Dataset (BUSI) [1]: This dataset contains breast ultrasound images collected from women between 25 and 75 years old.
-The dataset consists of 780 pairs of images in normal, benign, and malignant, with corresponding masks.
-Unlike the setting in [50], we include all three classes in the dataset.
-The dataset is split into the training set, the validation set and the validation set with a ratio of 8:1:1 with three different random states.
-4) CVC-ClinicDB  [4]:This dataset contains frames extracted from colonoscopy videos that contain several examples of polyps.
-The dataset consists of 612 pairs of images and corresponding masks.
-he dataset is split into the training set, the validation set and the testing set with a ratio of 8:1:1 with three different random states.
-5) CVC-ColonDB [3]: This dataset includes 380 images, each with a resolution of 574$\times$500 pixels.
-These images were selected from a pool of 12,000 images obtained from 15 short colonoscopy videos, with only these 380 images being annotated.
-The dataset is split into the training set, the validation set and the testing set with a ratio of 8:1:1 with three different random states.
-6) Kvasir-SEG  [25]: This dataset comprises 1000 polyp images and their corresponding ground truth masks with resolutions ranging from 332$\times$487 to 1920$\times$1072 pixels.
-he dataset is split into the training set, the validation set and the testing set with a ratio of 8:1:1 with three different random states.
-7) ISIC 2017  [13]: This dataset contains skin lesions images for Skin Lesion Analysis toward Melanoma Detection Challenge at the International Symposium on Biomedical Imaging (ISBI) 2017.
-There are 2000 pairs of images and corresponding masks in the training set, 150 pairs in the validation set, and 600 pairs in the testing set.
-8) GlaS  [49]: This dataset is designed for gland segmentation.
-It consists of 165 whole-slide images (WSIs), with 85 images designated for training and 80 for testing.
-The training dataset is split into a new training set and a validation set with a ratio of 9:1 with three different random states.
-Appendix B More Ablation and Exploration Analysis
-B.1 Depth-wise Convolution in IR-RWKV Blocks
-The comparison results to explore the effect of DW-Conv after Spatial Mix in IR-RWKV blocks are shown in Table A1.
-The inclusion of DW-Conv demonstrates a notable improvement in performance metrics.
-The slight increase in computational cost is minimal and well-justified given the significant gains in performance.
-This indicates that in IR-RWKV blocks, DW-Conv plays a critical role.
-It provides local information that complements the long-range dependencies captured by Spatial Mix, effectively enhancing model accuracy.
-Without pre-training, relying solely on Spatial Mix without the inclusion of local information from DW-Conv limits the model’s ability to fully capture fine-grained spatial details.
-This demonstrates the necessity of integrating both components to achieve a balanced representation of local and global features.
-Table A1: Comparison of results on the Synapse dataset: effect of applying DW-Conv after Spatial Mix in IR-RWKV blocks
-DW-Conv Average FLOPs
+0 w 12.72 69.83 09.18 G
+1 w/o 10.83 78.27 09.33 G
+1 w/ 10.37 83.19 10.96 G
+2 w/o 11.32 80.11 09.41 G
+2 w/ 11.62 84.28 11.04 G
+3 w/o 12.56 82.40 09.48 G
+3 (ours) w/ (ours) 08.85 84.29 11.11 G
+TABLE XIII: Comparison of different hidden rates in the ChannelMix of CCM on the Synapse dataset.
+Hidden Rate Average FLOPs
 HD95 $\downarrow$ DSC $\uparrow$
-w/o 16.55 75.56 11.14G
-w/ (ours) 14.03 78.68 11.22G
-B.2 Larger Resolution Inputs
-The comparison results of different methods with 512$\times$ 512 input in the Synapse dataset are shown in Table A2.
+1 10.47 82.94 10.57 G
+2 (ours) 8.85 84.29 11.11 G
+3 9.89 84.18 11.62 G
+4 9.24 84.06 12.14 G
+Based on the results in Table XIII, the hidden rate of the Channel Mix layer in the CCM module materially affects performance: increasing the rate from 1 to 2 yields a notable improvement in performance, while further increases bring no gains at the cost of higher FLOPs.
+Larger Resolution Inputs
+The comparison results of different methods with $\times 512$ input in the Synapse dataset are shown in Table XIV.
 The average Dice scores of other methods are from [73].
-Our results indicate that when the input resolution is increased to 512$\times$512, our model demonstrates an improvement in performance.
-Although the improvement is not significant, it still maintains the best performance overall.
-Furthermore, compared to TransUNet, the computational overhead associated with the resolution increase is less severe, showcasing the efficiency of our model at higher resolutions.
-Figure A1: #FLOPs vs input resolutions. for different methods. The input channel is 1 and the output is 9 (aligned with settings on Synapse dataset).
-Table A2: Comparison of different methods with 512$\times$512 resolution input on the Synapse dataset. The evaluation metric are DSC in (%). The average DSC of other methods are from  [10].
+Our results indicate that increasing the input resolution improves the performance of our model.
+Furthermore, compared to TransUNet, the computational overhead associated with the resolution increase is less severe, showcasing the efficiency of at higher resolutions.
+TABLE XIV: Comparison of different methods with $\times512$ resolution input on the Synapse dataset. The evaluation metric is DSC in (%). The average DSCs of CNN-based methods are from [73].
 Method Average FLOPs
@@
 nnU-Net [24] 82.92 -
-TransUNet  [9] 84.36 168.7G
-RWKV-UNet (ours)  [9] 85.52 058.6G
-As shown in Figure A1, the RWKV-UNet series demonstrates advantages in handling high-resolution inputs, as shown in the FLOPs comparison.
-Particularly for larger resolutions, the RWKV-UNet models maintain significantly lower computational costs compared to other high-performance architectures like transUNet and MT-UNet.
-It demonstrates that linear attention outperforms traditional Transformers in efficiency, enabling faster and more memory-efficient processing of high-resolution inputs and long sequences.
+TransUNet [28] 84.36 148.29G
+SAMed_h [76] 84.30 783.98G
+RWKV-UNet (ours) 86.73 058.05G
+V Conclusion
+In this study, we introduce RWKV-UNet, a novel architecture that integrates the RWKV architecture with U-Net.
+By combining the strengths of convolutional networks for local feature extraction and RWKV’s ability to model global context, our model significantly improves medical image segmentation accuracy.
+The proposed enhancements, including the GLSP module and the CCM module, contribute to a more precise representation of features and information fusion across different scales.
+Experimental results on 11 datasets demonstrate that RWKV-UNet surpasses SOTA methods.
+Its variants (RWKV-UNet-S and RWKV-UNet-T) offer a practical balance between performance and computational efficiency.
+Our approach has a strong potential to advance medical image analysis, particularly in clinical settings where both accuracy and efficiency are paramount.
+Limitations and Future Work.
+RWKV-UNet is a powerful 2D medical image segmentation model that effectively combines RWKV with convolutional operations; however, it is currently not applicable to 3D imaging.
+In the future, we plan to extend the model to 3D for handling volume data and to explore the potential of RWKV-based foundational models for medical imaging.
+We also aim to develop ultra-lightweight RWKV-based models tailored for point-of-care applications, preserving segmentation accuracy while enhancing adaptability and speed further.